### v1

# The Mechanics of Nucleation and Growth and the Surface Tensions of Active Matter

Abstract
Homogeneous nucleation, a textbook transition path for phase transitions, is typically understood on thermodynamic grounds through the prism of classical nucleation theory. However, recent studies have suggested the applicability of classical nucleation theory to systems far from equilibrium. In this Article, we formulate a purely mechanical perspective of homogeneous nucleation and growth, elucidating the criteria for the properties of a critical nucleus without appealing to equilibrium notions. Applying this theory to active fluids undergoing motility-induced phase separation, we find that nucleation proceeds in a qualitatively similar fashion to equilibrium systems, with concepts such as the Gibbs-Thomson effect and nucleation barriers remaining valid. We further demonstrate that the recovery of such concepts allows us to extend classical theories of nucleation rates and coarsening dynamics to active systems upon using the mechanically-derived definitions of the nucleation barrier and surface tensions. Three distinct surface tensions – the mechanical, capillary, and Ostwald tensions – play a central role in our theory. While these three surface tensions are identical in equilibrium, our work highlights the distinctive role of each tension in the stability of active interfaces and the nucleation and growth of motility-induced phases.
I Introduction
In recent years it has become clear that many of the intriguing patterns and phases formed by active systems do not occur spontaneously but require a rate-determining rare event [1, 2, 3, 4]. The rare events needed to initiate a variety of active phase transitions often appear remarkably similar to those of the corresponding process in passive systems. Phase separation and crystallization of homogeneous active fluids, under certain conditions, require the formation of a spherical critical nucleus [2, 5]. Upon forming, the nuclei can continue to grow and coarsen through the familiar coarsening mechanisms of coalescence and Ostwald ripening [6, 7, 8]. These observations suggest that classical nucleation theory (CNT) and established coarsening theories may be applicable to active systems [9, 10, 11, 12, 13, 14, 15]. CNT has shaped and framed our understanding of the transition rate of a variety of activated processes in passive systems including the rupture of biological membranes [16], the formation of dust clouds on exoplanets [17], and the motion of ferroelectric domain walls [18]. In its canonical form, CNT implicitly leverages microscopic time-reversal symmetry (TRS) to posit a free energy barrier for the rare event. In the case of a metastable fluid that will ultimately crystallize or phase separate, the rare event required to initiate the phase transition is the formation of a spherical finite-sized nucleus. Straightforward thermodynamic arguments weighing the competing effects of the volumetric driving force for forming a nucleus and the associated interfacial penalty allow for the determination of the barrier height and the critical nucleus’ density and radius using bulk equations of state (e.g., the chemical potential) and the surface energy.
The naive application of the above ideas to active systems results in immediate issues: concepts such as chemical potential [19, 20] and surface energy [21, 22, 23, 24, 25, 26] are not uniquely defined in active systems. Nevertheless, a number of simulations and numerical studies in the last decade have made it clear that CNT and post-nucleation coarsening theories may be applicable far from equilibrium. Zakine et al. developed [27] and applied [28] action minimizing methods to stochastic field dynamics described by Active Model B [29] and found that the transition from homogeneous to inhomogeneous states is well described by the rate-limiting formation of a radially symmetric nucleus. Particle-based simulations of metastable active fluids undergoing motility-induced phase separation (MIPS) conducted by Redner et al. [2] were consistent with the cluster rate formation statistics posited by Becker-Döring theory [9, 12, 15]. These observations suggest that one may be able to construct a theory for the nucleation of active phases with similar physical assumptions as CNT – but without invoking a thermodynamic variational principle.
Significant progress towards generalizing CNT to nonequilibrium fluids has been recently made by Cates, Nardini, and co-workers [23, 4]. Introducing the minimal stochastic field theory, Active Model B+ (AMB+), they demonstrated that the size and density of the critical nucleus can be solely determined from dynamics [23] and elucidated the conditions at which the critical nucleus represents an unstable fixed point. Furthermore, they derived the radial growth dynamics of active nuclei, finding theoretical justification for an effective free energy landscape that controls the size distribution of pre-critical nuclei [4]. Although these efforts have yielded crucial insights towards the nucleation of active phases, explicitly mapping microscopic models of active systems to the phenomenological coefficients of AMB+ may not always be possible. Moreover, the absence of a unique mechanical interpretation of active field theories makes it difficult to understand the physical significance of the negative mechanical surface tension [21] (which does control the Laplace pressure difference of active bubbles and droplets [30]) of active particles in nucleation.
In this Article, we seek to answer the following questions beginning from a microscopic description and without appealing to equilibrium notions: Provided the saturation conditions, how might one determine the size ($R_{C}$) and density ($\rho^{\rm in}$) of the critical nucleus?; what is the precise role of the various active surface tensions in the nucleation process?; and how might one describe the post-nucleation growth dynamics of coarsening active fluids? To address these questions, we begin by deriving the conditions for determining the density and radius of a critical nucleus from purely mechanical considerations. We apply this framework to fluid nucleation of active Brownian particles (ABPs) and determine the critical nucleus properties as a function of activity and the saturation density. We then combine the recently developed perspective of Cates and Nardini [4] with our fluctuating hydrodynamics of ABPs [26] to describe the time evolution of the critical nucleus size, finding that droplets of ABPs are associated with a far higher nucleation barrier than bubbles. This nucleation barrier follows the familiar equilibrium intuition of an energetic penalty scaling with the surface area of the nucleus competing with an energetic driving force scaling with the volume of the nucleus. Building off our active CNT, we are able to extend classic perspectives on coarsening, including those of Lifhitz, Sylozov, and Wagner [11] as well as Voorhees and Glicksman [31, 32], to systems far from equilibrium. We are thus able to numerically investigate large-scale, late-stage coarsening dynamics that would otherwise be prohibitive for particle-based simulations.
Surprisingly, we find that the properties and dynamics of the critical nucleus of ABPs are dependent on three different surface tensions. These tensions include the negative mechanical surface tension [21], which controls the pressure difference between the nucleus and parent phase, the capillary-wave tension [25, 26], which drives interfaces to minimize surface area and results in a spherical nucleus geometry, and the Ostwald tension [23], which plays a central role in both the nucleation barrier and coarsening dynamics. At equilibrium, all three of these tensions must be identical and are defined by the energetic cost of increasing interfacial area, but far from equilibrium they are free to take on disparate values. Despite this nonequilibrium peculiarity, we demonstrate the recovery of phenomena typically associated with equilibrium, such as the Gibbs-Thomson effect and nucleation barriers, from dynamics that microscopically violate time reversal symmetry.
II Results and Discussion
II.1 Equilibrium Nucleation Theory
We briefly recapitulate the thermodynamic arguments of CNT applied to bubble/droplet nucleation, the focus of this work. The complete derivation can be found in Section 3 of the Supporting Information (SI)  [^1] See Supporting Information at [URL]. We consider a macroscopic homogeneous single-component system prepared with a uniform number density $\rho^{\rm sat}$ within a fixed system volume $V$ (the total number of particles is thus also fixed at $N=V\rho^{\rm sat}$). We focus on saturation densities $\rho^{\rm sat}$ within the liquid-gas binodal of the phase diagram but far from the spinodal boundaries, limiting our analysis to be far from the critical point. In this region of the phase diagram, a system prepared with a spatially uniform density of $\rho^{\rm sat}$ will remain linearly stable against density fluctuations until a critical nucleus with density $\rho^{\rm in}$ and volume $V^{\rm nuc}$ is formed. We assume that the formation of the critical nucleus is a sufficiently rare event that we may consider a single nucleation event in isolation and a nucleus volume, $V^{\rm nuc}$, that is much smaller than the system volume. These assumptions provide additional limits on the range of $\rho^{\rm sat}$ in which CNT is applicable.
The extensive Helmholtz free energy for the system is simply $\mathcal{F}=V^{\rm nuc}f(\rho^{\rm in})+\left(V-V^{\rm nuc}\right)f(\rho^{\rm out})+\partial V^{\rm nuc}\upgamma$ where $f(\rho)$ is the free energy density of a uniform fluid with density $\rho$, the density of the parent phase surrounding the nucleus is $\rho^{\rm out}$, and $\upgamma(>0)$ is the surface energy. The interfacial penalty ensures that the mean-field nucleus shape will be spherical in order to reduce surface area, which we have labeled as $\partial V^{\rm nuc}$. Conservation of the total particle number results in a constraint of $\rho^{\rm in}V^{\rm nuc}+\rho^{\rm out}\left(V-V^{\rm nuc}\right)=\rho^{\rm sat}V$. Note that while the spatial density profile must smoothly vary from the nucleus density $\rho^{\rm in}$ to the parent phase density $\rho^{\rm out}$ this interfacial width is much smaller than the nucleus size and its contribution to the particle conservation constraint is negligible. This allows us to focus on the coexisting interior and exterior densities rather than the complete spatial profile. We then extremize the free energy (subject to the particle conservation constraint) with respect to the coexisting densities and nucleus size and find the following conditions:
$\mu(\rho^{\rm in})$ $=\mu(\rho^{\rm out}),$ (1a)
$p(\rho^{\rm in})$ $=p(\rho^{\rm out})+\frac{2\upgamma}{R_{C}},$ (1b)
where we have defined the chemical potential $\mu(\rho)\equiv\partial f/\partial\rho$ and pressure $p(\rho)\equiv-f(\rho)+\rho\mu(\rho)$. Here, $R_{C}$ is the critical radius of the spherical nucleus. These criteria are the expected conditions of chemical potential equality and the Young-Laplace pressure difference which, due to the positive surface energy, results in a higher nucleus pressure relative to the parent phase. Together with the particle conservation constraint, Eqs. (1) can be used to solve for $\rho^{\rm in}$, $\rho^{\rm out}$ and $R_{C}$.
We can make use of the condition that the droplet volume is much smaller than the system volume such that $\rho^{\rm out}\approx\rho^{\rm sat}$. In this limit, we have effectively changed our thermodynamic ensemble as the nucleus is now in contact with a particle reservoir with fixed chemical potential $\mu(\rho^{\rm sat})\equiv\mu^{\rm sat}$ and thus $\rho^{\rm in}$ can be found directly from $\mu(\rho^{\rm in})=\mu^{\rm sat}$. The critical radius is then simply ${R_{C}=2\upgamma/(\mu^{\rm sat}\Delta\rho-\Delta f})$ where the difference between an arbitrary property, $a$, in the interior and exterior of the nucleus is defined as $\Delta a\equiv a^{\rm in}-a^{\rm sat}$. The free energy barrier associated with the critical nucleus can be expressed as the sum of the volumetric and interfacial contributions ${F^{\rm{\dagger}}=4\pi R_{C}^{2}\upgamma-\frac{4}{3}\pi R_{C}^{3}(\mu^{\rm sat}\Delta\rho-\Delta f)}$ or equivalently, ${F^{\rm{\dagger}}=\frac{4}{3}\pi\upgamma R_{C}^{2}}$ [33]. The probability of observing such a critical nucleus is thus expected to be $P(R_{C})\sim\text{exp}\left[F^{{\dagger}}/k_{B}T\right]$.
It is clear that in passive systems, thermodynamics alone provides a guide for determining the density, size, and statistics of the critical nucleus. Given a saturation density, Eqs. (1) are the two equations that allow for the determination of the two unknowns: $\rho^{\rm in}$ and $R_{C}$. In the following section, we derive analogous equations capable of determining $\rho^{\rm in}$ and $R_{C}$ for systems arbitrarily far from equilibrium by adopting a mechanical perspective.
II.2 Mechanics of a Critical Nucleus
The pathway for transitioning from a homogeneous supersaturated active fluid to a state of two-fluid coexistence appears eerily reminiscent to those of passive systems [2, 28]. In both scenarios, the initial activated process is the formation of a spherical nucleus with a specific critical radius $R_{C}$ and internal density $\rho^{\rm in}$. In the absence of a free energy, the traditional arguments previously outlined used to establish passive classical nucleation theory are no longer applicable. We therefore take a generalized mechanical approach with the aim of extending CNT to systems arbitrarily far from equilibrium. We again limit our analysis to saturation densities near the liquid and gas binodal densities such that the density of the fluid surrounding a critical nucleus is approximately $\rho^{\rm sat}$ and we may focus on a single nucleus coexisting with this saturated fluid.
Let us consider the dynamical definition of a critical nucleus. When the critical nucleus is an unstable fixed point  [^2] Recently Tjhung et al. [23] demonstrated in some active systems the critical nucleus may be a stable fixed point, in such a case nuclei smaller (larger) than $R_{C}$ will grow (shrink)., nuclei smaller (larger) than $R_{C}$ will be associated with a particle flux normal to the interface such that the nuclei shrink (grow). $R_{C}$ thus represents a critical size at which the direction of this flux switches, so the system must have zero flux normal to the interface precisely when a nucleus has a radius of $R_{C}$. For a single realization of an active fluid, the stochastic noise plays a crucial role in the density flux, which we will discuss further in the next section. For the purposes of determining the properties of the critical nucleus, we can consider an ensemble of supersaturated states. We choose a spherical coordinate system with an origin coinciding with that of the nucleus. The ensemble-averaged nucleus shape will be spherical and the ensemble-averaged density flux will no longer be stochastic. We will determine the critical nucleus size by searching for states of zero radial flux.
Two flux-free states are well-established: the macroscopically phase separated state that the system will eventually reach and the initial metastable homogeneous state. The critical nucleus will correspond to an additional state in which a finite-size minority phase with zero radial flux coexists with a parent phase. The condition for determining a stationary state of a density profile is given by the linear momentum balance (see Section 1 of the SI [33] and Ref. [35]) which governs the dynamics of the density flux. On timescales where inertia can be neglected (or for all times for overdamped particle dynamics), the radial-component of the linear momentum balance is simply $\left(\bm{\nabla}\cdot\bm{\sigma}+\mathbf{b}\right)\cdot\mathbf{e}_{r}=0$, where $\bm{\sigma}$ is the stress tensor, $\mathbf{b}$ are the body forces, and $\mathbf{e}_{r}$ is the unit vector in the radial dimension. We note that exact microscopic expressions for both $\bm{\sigma}$ and $\mathbf{b}$ can be determined through an Irving-Kirkwood-Noll procedure [36, 37, 38], an approach that we will later adopt in applying this theory to active Brownian particles. In these formulations, terms often appear to be non-local, requiring integrals over all space to, for instance, compute the stress at a particular point in space arising from particle interactions. Nevertheless, these terms are often well-approximated (e.g., when interactions are short-ranged and partial spatial correlations decay quickly) by a series of local terms, the coefficients of which can be systematically determined [39]. We thus focus on systems in which $\bm{\sigma}$ and $\mathbf{b}$ can be fully expressed as local density gradient expansions.
Up to a divergence-free term, we may absorb the body force into a stress-like contribution such that ${\bm{\nabla}\cdot\bm{\sigma}+\mathbf{b}=\bm{\nabla}\cdot\left[\bm{\sigma}+\mathbf{F}\right]=\bm{\nabla}\cdot\bm{\Sigma}}$, where $\mathbf{F}$ is the Coulomb integral of $\mathbf{b}$ and $\bm{\Sigma}\equiv\bm{\sigma}+\mathbf{F}$ is the dynamic stress tensor. While $\bm{\sigma}$ and $\mathbf{b}$ are fully local, the Coulomb integral may result in non-local contributions to $\mathbf{F}$. The momentum balance (${\bm{\nabla}\cdot\bm{\Sigma}=\bm{0}}$) can be used to solve for the density profile upon identifying the precise form of the dynamic stress tensor. Under spherical symmetry, $\bm{\Sigma}$ has only two independent components and can be expressed as $\bm{\Sigma}=\Sigma_{rr}\mathbf{e}_{r}\mathbf{e}_{r}+\Sigma_{tt}\left(\mathbf{I}-\mathbf{e}_{r}\mathbf{e}_{r}\right)$ where $\mathbf{I}$ is the identity tensor. We can make analytical progress by expanding the dynamic stress tensor to second order in density gradients while leaving any potential non-local contributions symbolic:
$\Sigma_{rr}=$ $-\mathcal{P}(\rho)+\kappa_{1}(\rho)\left(\frac{\partial}{\partial r}+\frac{2}{r}\right)\frac{\partial\rho}{\partial r}$
$+\bigl{(}\kappa_{2}(\rho)+\kappa_{3}(\rho)\bigr{)}\left(\frac{\partial\rho}{\partial r}\right)^{2}+\kappa_{4}(\rho)\frac{\partial^{2}\rho}{\partial r^{2}}+\mathcal{K}_{r},$ (2a)
$\Sigma_{tt}=-\mathcal{P}(\rho)+\kappa_{1}(\rho)\left(\frac{\partial}{\partial r}+\frac{2}{r}\right)\frac{\partial\rho}{\partial r}+\kappa_{3}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\mathcal{K}_{t},$ (2b)
where $-\mathcal{P}(\rho)$ is the bulk isotropic contribution to the dynamic stress tensor, $\mathcal{K}_{r/t}$ are the non-local terms originating from the Coulomb integral of $\mathbf{b}$, and $\{\kappa_{i}\}$ are interfacial coefficients. These interfacial coefficients can be found by formally carrying out the expansion of $\bm{\sigma}$ and $\mathbf{b}$ beginning from the microscopic definition of stress and expanding the local terms that arise from the Coulomb integral of $\mathbf{b}$. The procedure for determining the precise forms of $\{\kappa_{i}\}$ and $\mathcal{K}_{r/t}$ from the expansion coefficients of $\bm{\sigma}$ and $\mathbf{b}$ is outlined in Section 1 of the SI [33]. In a passive system $\mathcal{P}(\rho)$ is simply the bulk pressure and $\{\kappa_{i}\}$ are given by the Korteweg expansion [40] while non-local terms can be well-approximated by local terms for single-component systems with short-ranged interactions. More generally, the bulk pressure and interfacial coefficients can contain dynamical contributions otherwise absent in passive systems [26].
Given the above physical picture, we now outline our process to find the zero radial flux states associated with a finite-sized minority phase [33]. Our spherically symmetric geometry results in the linear momentum balance simplifying to:
$0=\frac{\partial\Sigma_{rr}}{\partial r}-\frac{2}{r}\mathcal{G}(\rho),$ (3)
where we have defined $\mathcal{G}(\rho)=\Sigma_{tt}-\Sigma_{rr}$ as the “excess stress” in the non-radial directions. Deep within the both the nucleus and parent phase, where density gradients vanish, the dynamic stress tensor must be isotropic, implying $\mathcal{G}(\rho)$ is zero everywhere except within the interfacial region. If we assume that the width of the interfacial region, $w$, is small compared to the radius of the critical nucleus, integration of Eq. (3) from $r=0$ to $r=\infty$ results in:
$\Delta\mathcal{P}=\Delta\mathcal{K}_{r}+\frac{2\upgamma_{\rm mech}}{R_{C}}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right).$ (4a)
Here, $\upgamma_{\rm mech}$ is the mechanical surface tension as defined by Kirkwood and Buff [41]:
$\upgamma_{\rm mech}$ $=\int_{0}^{\infty}\left[\Sigma_{tt}-\Sigma_{rr}\right]\text{d}r,$
$=-\int_{0}^{\infty}\left[\kappa_{3}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\kappa_{4}(\rho)\frac{\partial^{2}\rho}{\partial r^{2}}+\mathcal{K}_{r}-\mathcal{K}_{t}\right]\text{d}r,$ (4b)
where the second equality utilizes our specific form of the dynamic stress [see Eqs. (2)]. Equation (4a) is a generalized Young-Laplace equation – with the equilibrium surface energy and pressure more generally replaced by the mechanical tension and dynamic pressure, respectively, and an additional contribution arising from the non-local terms. This mechanical tension was determined to be strongly negative for phase-separated ABPs [21] and has been the subject of some controversy regarding its physical relevance. Solon et al. [30] correctly identified that the mechanical surface tension of Kirkwood and Buff controls the Laplace pressure difference of droplets consistent with a fully local (dynamic) stress. More generally, $\Delta\mathcal{K}_{r}$ may also contribute to the Laplace pressure difference.
While the generalized Young-Laplace equation provides a mechanical condition remarkably similar to its equilibrium counterpart, we still require a second condition to allow us to determine our two unknowns, $\rho^{\rm in}$ and $R_{C}$. For passive fluids, this condition is equality of chemical potential between the nucleus and the surrounding phase. The search for chemical potential-like quantities for nonequilibrium systems has been the subject of much interest in the last decade in the context of determining macroscopic binodals [19, 42, 20, 35]. These works have resulted in generalized Maxwell constructions [20, 35] that are derived from performing the appropriate integration of the mechanical criterion of coexistence [43, 35].
We follow similar arguments used in constructing the nonequilibrium coexistence criteria of macroscopic binodals [43, 35]. Beginning from the mechanical balance [Eq. (3)] we perform an indefinite integral to find an integro-differential equation for $\Sigma_{rr}$ with the Laplace pressure used as a boundary condition. We then introduce the generalized Maxwell construction pseudovariable [43, 20, 35] $\mathcal{E}$ defined by the differential equation ${(\kappa_{1}+\kappa_{4})\partial^{2}\mathcal{E}/\partial\rho^{2}=(2(\kappa_{2}+\kappa_{3})-\partial(\kappa_{1}+\kappa_{4})/\partial\rho)\partial\mathcal{E}/\partial\rho}$, which is found to be $\mathcal{E}\sim 1/\rho$ for a passive system. We multiply the indefinite integral of Eq. (3) by $d\mathcal{E}/dr$ and integrate from $r=0$ to $r=\infty$ to arrive at:
$\int_{\mathcal{E}^{\rm in}}^{\mathcal{E}^{\rm sat}}\text{d}\mathcal{E}$ $\left[\mathcal{P}(\rho)-\mathcal{P}(\rho^{\rm sat})\right]=\frac{2\upgamma_{\rm ost}}{R_{C}}\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right),$ (5a)
where, analogously to Tjhung et al. [23], we have defined the “Ostwald” tension as
$\upgamma_{\rm ost}=\frac{\Delta\rho}{\rho^{\rm sat}\Delta\mathcal{E}}$ $\Biggl{[}\int_{0}^{\infty}\Bigl{(}\kappa_{1}(\rho)\frac{\partial\mathcal{E}}{\partial\rho}\left(\frac{\partial\rho}{\partial r}\right)^{2}$
$+\mathcal{G}$ $\left(\mathcal{E}-\mathcal{E}^{\rm in}\right)+\frac{R_{C}}{2}\frac{\partial\mathcal{K}_{r}}{\partial r}\left(\mathcal{E}^{\rm in}-\mathcal{E}\right)\Bigr{)}\text{d}r\Biggr{]}.$ (5b)
Note that $\upgamma_{\rm ost}$ has been defined such that it has the same units as mechanical surface tension and becomes identical to $\upgamma_{\rm mech}$ in a passive system [33].
The connection between Eqs. (4a) and (5a) to the thermodynamic criteria shown in Eqs. (1) can be made explicit. As described by Solon et al., the generalized Maxwell constructions on the pressure can also be expressed as a difference in chemical potential-like quantities between coexisting phases [30]. It was subsequently shown that this potential satisfies a generalized Gibbs-Duhem relation [44], which, for the single-component system under consideration, is expressed as $\text{d}g=\mathcal{E}\text{d}\mathcal{P}$ where $g$ is the generalized chemical potential. The generalized Gibbs-Duhem relation allows [33] Eq. (5a) to be expressed as a difference of chemical potentials with the nonequilibrium critical nucleus criteria now taking the form:
$\Delta g=\mathcal{E}^{\rm in}\Delta\mathcal{K}_{r}+\frac{2}{R_{C}}\left[\mathcal{E}^{\rm in}\upgamma_{\rm mech}-\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}\upgamma_{\rm ost}\right],$ (6a)
$\Delta\mathcal{P}=\Delta\mathcal{K}_{r}+\frac{2\upgamma_{\rm mech}}{R_{C}}.$ (6b)
Additionally, by defining a pseudopotential ${\Phi\equiv\int\mathcal{P}\text{d}\mathcal{E}}$, Eq. (5a) can be rearranged to express the critical radius as ${R_{C}=2\upgamma_{\rm ost}\rho^{\rm sat}\Delta\mathcal{E}/\Delta\rho\left(\Delta\mathcal{E}\mathcal{P}^{\rm sat}-\Delta\Phi\right)}$. For passive systems without non-local terms, $\upgamma_{\rm mech}=\upgamma_{\rm ost}$ and $\mathcal{E}\sim 1/\rho$. One can use these relations to readily verify that $g^{\rm in}=g^{\rm sat}$ and that Eq. (5a) recovers the thermodynamically derived expression for $R_{C}$ [33]. Our mechanical criteria for the critical nucleus [Eqs. (6)] thus recover and generalize the thermodynamic criteria [Eq. (1)].
Equations (6) [or, equivalently, Eqs. (4) and (5)] are the finite-size coexistence criteria for the critical nucleus. In Ref. [23], similar finite-sized coexistence criteria were derived with two key differences. First, the Laplace pressure difference in Ref. [23] was found to be controlled by the tension controlling Ostwald ripening rather than $\upgamma_{\rm mech}$. Additionally, the chemical potentials (including non-local terms) inside and outside of the nucleus were found to be equal by Ref. [23]. Fundamentally, this is because pressure and chemical potential have different physical interpretations in Ref. [23] than in this work. Here, pressure is the bulk isotropic component of the dynamic stress tensor and carries the familiar mechanical interpretation, while in Ref. [23] the pressure (also referred to as a “pseudopressure”) cannot be physically interpreted without identifying the pseudovariable. In Ref. [23] the chemical potential is defined as the scalar field controlling the dynamics and carries much of the familiar intuition of a chemical potential, while $g$ in this work cannot be physically interpreted without identifying $\mathcal{E}$. The mechanical approach presented above may be a more natural starting point when beginning with microscopic equations-of-motion and using an Irving-Kirkwood approach to derive the form of the dynamic stress and offers the advantage of clearly identifying the role of the mechanical tension of Kirkwood and Buff.
Given a saturation density, one may use Eqs. (6) to solve for the unknowns $\rho^{\rm in}$ and $R_{C}$. We note that as $\rho^{\rm sat}$ approaches either binodal density, $R_{C}$ will diverge; Eq. (4a) and Eq. (5a) then simplify to the macroscopic phase coexistence criteria presented in Ref. [35]. As one approaches this limit and $R_{C}$ grows, $\mathcal{O}\left(R_{C}^{-2}\right)$ corrections to the coexistence criteria become negligible and we expect our theory to be most accurate. Solution of the coexistence criteria is complicated by the presence of $\upgamma_{\rm mech}$ and $\upgamma_{\rm ost}$ in Eqs. (6), which depend on the density profile, which in turn depends on $R_{C}$ and $\rho^{\rm in}$. Therefore, solution of Eqs. (6) requires self-consistently solving for the density profile using a procedure detailed in Section 1 of the SI [33]. In the following section, we apply our mechanical theory to a canonical model system for active matter, active Brownian particles.
II.3 Nucleation of Motility-Induced Phases
We consider a system of $N$ interacting ABPs in which the time variation of the position $\mathbf{r}_{i}$ and orientation $\mathbf{q}_{i}$ of the $i$th particle follow overdamped Langevin equations:
$\dot{\mathbf{r}}_{i}$ $=U_{o}\mathbf{q}_{i}+\frac{1}{\zeta}\sum_{j\neq i}^{N}\mathbf{F}_{ij},$ (7a)
$\dot{\mathbf{q}}_{i}$ $=\mathbf{q}_{i}\times\bm{\Omega}_{i},$ (7b)
where $\mathbf{F}_{ij}$ is the (pairwise) interparticle force, $U_{o}$ is the intrinsic active speed, $\zeta$ is the translational drag coefficient, and $\bm{\Omega}_{i}$ is a stochastic angular velocity with zero mean and variance $\langle\bm{\Omega}_{i}(t)\bm{\Omega}_{j}(t^{\prime})\rangle=2D_{R}\delta_{ij}\delta(t-t^{\prime})\mathbf{I}$. Here, $D_{R}$ is the rotational diffusivity, which may be athermal in origin, and is inversely related to the orientational relaxation time $\tau_{R}\equiv D_{R}^{-1}$. For the remainder of this Article, we consider purely repulsive $\mathbf{F}_{ij}$, corresponding to a sufficiently stiff interaction potential such that hard-sphere statistics are closely approximated with an effective hard-sphere diameter of $d_{\rm hs}$ [45, 5]. One can then define an intrinsic run length, $\ell_{o}/d_{\rm hs}=U_{o}\tau_{R}$, the typical distance an ideal particle travels before reorienting.
Figure 1: Density profiles of critical bubbles and droplets at $\ell_{o}/d_{\rm hs}=150$. All profiles have been shifted by their corresponding critical radius, given in Fig. 2. Here, the volume fraction is defined as the number density multiplied by the effective volume of a particle $\rho\pi d_{\rm hs}^{3}/6$.
The form of $\mathcal{P}(\rho)$ and $\{\kappa_{i}(\rho)\}$ for ABPs are given by [35, 26]:
$\mathcal{P}(\rho)$ $=p_{C}(\rho)+\frac{\rho\zeta U_{o}\ell_{o}\overline{U}(\rho)}{d(d-1)},$ (8a)
$\kappa_{1}(\rho)$ $=\frac{3\ell_{o}^{2}}{2d(d-1)(d+2)}\left(\overline{U}(\rho)\right)^{2}\frac{\partial p_{C}}{\partial\rho},$ (8b)
$\kappa_{3}(\rho)$ $=\frac{3\ell_{o}^{2}}{2d(d-1)(d+2)}\overline{U}(\rho)\frac{\partial}{\partial\rho}\left[\overline{U}(\rho)\frac{\partial p_{C}}{\partial\rho}\right],$ (8c)
$\kappa_{2}(\rho)$ $=\kappa_{4}(\rho)=0,$ (8d)
$\mathcal{K}_{r}$ $=\mathcal{K}_{t}=0,$ (8e)
where $p_{C}(\rho)$ is the conservative pressure arising from particle interactions, $\overline{U}(\rho)$ is the dimensionless effective active speed, and $d(>1)$ is the system dimensionality. In Eqs. (8) we have ignored the gradient contributions of the interaction potential to $\bm{\Sigma}$, which is justified when $\ell_{o}/d_{\rm hs}>>1$ for purely repulsive interactions [30, 35, 44]. Given our expressions for $\{\kappa_{i}(\rho)\}$, one can solve for the generalized Maxwell construction pseudovariable $\mathcal{E}\sim p_{C}(\rho)$ [35]. We also note that because $\mathcal{K}_{r}=0$, Eq. (4a) reduces to the familiar Laplace pressure difference (i.e. no non-local corrections) and the negative sign of $\upgamma_{\rm mech}$ [21] implies that the pressure inside of a droplet is lower than the pressure outside of the droplet [30]. The phase diagram, dependent on $\ell_{o}/d_{\rm hs}$ and volume fraction $\phi\equiv\pi d_{\rm hs}^{3}/6$, has been well established computationally [5] and theoretically [35] with the latter predicting a critical point at $\phi\approx 0.49$ and $\ell_{o}\approx 16.2d_{\rm hs}$
Figure 2: Critical radius and relative change in nucleus density (${\nu^{\rm in}\equiv|\rho^{\rm in}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$) as a function of supersaturation (${\nu^{\rm sat}\equiv|\rho^{\rm sat}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$).
With the dynamic stress of ABPs in hand, we are able to self-consistently solve our generalized critical nucleus conditions [Eqs. (6)], with the resulting density profiles for both bubbles and droplets shown in Fig. 1 for a select activity and various supersaturations. Values of $R_{C}$ and $\rho^{\rm in}$ obtained from self-consistent solution of the criteria at various activities and supersaturations are provided in Fig. 2. Here, we define a dimensionless overall supersaturation density as as the relative departure of the majority phase density from the binodal density, ${\nu^{\rm sat}=|\rho^{\rm sat}-\rho^{\rm gas,liq}|/\rho^{\rm gas,liq}}$ and a dimensionless nucleus density as the relative departure of the minority phase density from the opposite binodal ${\nu^{\rm in}=|\rho^{\rm in}-\rho^{\rm liq,gas}|/\rho^{\rm liq,gas}}$. Our theory is expected to hold as $\nu^{\rm sat}\to 0$. We observe a roughly linear relationship between $\nu^{\rm sat}$ and $\nu^{\rm in}$ as well as an inverse relationship between $\nu^{\rm sat}$ and $R_{C}$. These results are therefore in excellent agreement with existing simulation data reporting that ABPs obey a Gibbs-Thomson relation with $\nu^{\rm in}\propto 1/R$ [22].
Figure 2 displays a striking difference between critical bubbles and droplets. At all examined values of activity, critical droplets of active Brownian particles are over an order of magnitude larger than critical bubbles under similar saturation conditions. This difference between bubbles and droplets widens as activity is increased and the phases become increasingly distinct. To contextualize this order of magnitude difference, we calculate the critical nucleus size and density for an equilibrium van der Waals fluid [33]. While the size of the critical droplets are indeed larger than the size of critical bubbles in passive fluids, they remain within an order of magnitude of each other at similar saturation conditions even far from the critical point. In equilibrium, this difference is rooted in the asymmetry of the equations of state outside of the binodal densities: the vapor bubbles are more compressible than the liquid droplets. This asymmetry results in larger Laplace pressure differences for bubbles in comparison to droplets under similar saturation conditions, leading to the critical radius discrepancy. While this asymmetry in the compressibility is also present for active systems, the coexistence criteria contain uniquely nonequilibrium contributions, including a significant discrepancy between the vapor and bubble Ostwald tensions. At equilibrium these two quantities are identical, but can differ by an order of magnitude in active systems as shown in Fig. 3.
II.4 Langevin Dynamics of a Nucleus
Further insight into the active nucleation process can be obtained by considering the growth dynamics of nuclei near their critical size. We assume the dynamics of the nucleus to proceed quasistatically: any given nucleus with radius $R$ (not necessarily equal to $R_{C}$) will be in a local mechanical balance with its surroundings. We now distinguish the density just outside the nucleus (now denoted as $\rho^{\rm out}$), and the density far from the nucleus which remains fixed at the saturation density, $\rho^{\rm sat}$. The mechanical balance between the nucleus and its immediate surroundings will determine the intermediate density which will generally not equal the far away saturation density, $\rho^{\rm out}\neq\rho^{\rm sat}$. The inequality between the intermediate and far away densities will drive the transport of particles and will cause $\rho^{\rm out}$ to evolve in time. This, in turn, will cause $R$ to also change in time as determined by the local mechanical balance. Cates and Nardini [4] recently demonstrated that such a description can allow for the determination of effective nucleation barriers of active fluids with dynamics described by AMB+, despite the absence of well-defined free energy for these driven systems. Deriving the stochastic growth dynamics of the nuclei requires a fluctuating hydrodynamic description of the density field which we recently obtained for interacting ABPs [26]:
$\frac{\partial\rho}{\partial t}$ $=-\bm{\nabla}\cdot\mathbf{J},$ (9a)
$\mathbf{J}$ $=\frac{1}{\zeta}\bm{\nabla}\cdot\bm{\Sigma}+\bm{\eta}^{\rm act},$ (9b)
where $\rho$ now is a stochastic density field (while in the previous sections it was an ensemble average), $\bm{\Sigma}$ is defined identically as the dynamic stress tensor of the previous section and $\bm{\eta^{\rm act}}$ is an athermal stochastic contribution to the flux with zero mean and a variance of:
$\langle\bm{\eta}^{\rm act}(\mathbf{r},t)\bm{\eta}^{\rm act}(\mathbf{r}^{\prime},t^{\prime})\rangle=$ $2\frac{k_{B}T^{\rm act}}{\zeta}\left(\rho\mathbf{I}-\frac{d}{d-1}\mathbf{Q}^{\prime}\right)$
×$\times\delta(t-t^{\prime})\delta(\mathbf{r}-\mathbf{r}^{\prime})\ .$ (9c)
Here $\mathbf{Q}^{\prime}$ is the traceless nematic order and we have defined $k_{B}T^{\rm act}=\ell_{o}\zeta U_{o}/d(d-1)$ as the active energy scale.
We recently established a surface-area minimizing principle for liquid-gas interfaces of ABPs resulting from a positive capillary-wave tension, $\upgamma_{\rm cw}$, which resists fluctuations to the area-minimizing interfacial shape. The strength of the fluctuations were found to be set by the athermal active energy scale $k_{B}T^{\rm act}$ [26]. The ratio of the capillary tension to this energy scale – the active interfacial stiffness – controls the amplitude of long-wavelength interfacial fluctuations. A fluctuating nucleus can be well-approximated as a sphere when the average radius $R$ satisfies $(R^{2}\upgamma_{\rm cw}/k_{B}T^{\rm act})^{1/2}>>1$. Under this condition, the stochastic density field dynamics can be connected to the radial growth dynamics of the nucleus through the following radially symmetric ansatz [46, 4]:
[…]
$\zeta_{\rm eff}\frac{\partial R}{\partial t}$ $=F^{D}(R)+\tilde{\chi}(t),$ (11a)
$\langle\tilde{\chi}(t)\rangle$ $=0,$ (11b)
$\langle\tilde{\chi}(t)\tilde{\chi}(t^{\prime})\rangle$ $=2k_{B}T^{\rm act}\zeta_{\rm eff}\delta(t-t^{\prime}),$ (11c)
$\zeta_{\rm eff}$ $=\frac{4\pi\Delta\rho^{2}R^{3}\zeta}{\rho^{\rm sat}},$ (11d)
$F^{D}(R)$ $=-8\pi\upgamma_{\rm ost}R\left[1-\frac{R}{R_{C}}\right].$ (11e)
The radial growth dynamics take the form of a one-dimensional overdamped Langevin equation containing three distinct forces: a drag force with drag coefficient $\zeta_{\rm eff}$, a stochastic force with an athermal temperature scale of $k_{B}T^{\rm act}$ [47], and a deterministic force $F^{D}(R)$ proportional to the Ostwald tension. This deterministic force vanishes when $R=R_{C}$ where the density outside nucleus is identical to the far away saturation density $\rho^{\rm out}=\rho^{\rm sat}$. Crucially, the noise statistics satisfy the fluctuation-dissipation theorem (FDT) despite the underlying microscopic dynamics [Eqs. (7)] violating FDT. We can plainly see that $\upgamma_{\rm ost}$ is the quantity controlling the dynamics of nucleation and ripening. As first argued by Thjung et al. [23] a positive $\upgamma_{\rm ost}$ implies $R_{C}$ is an unstable fixed point in which nuclei with $R>R_{C}$ tend to grow while nuclei with $R<R_{C}$ tend to shrink. A priori, there is no reason to assume $\upgamma_{\rm ost}$ is generally positive. In the case of a negative $\upgamma_{\rm ost}$, $R_{C}$ becomes a stable fixed point. However, for ABPs we do not find a negative $\upgamma_{\rm ost}$ [see Fig. 4] – implying that nucleation occurs with the same qualitative dynamics as an equilibrium system as observed in simulation [2].
Figure 3: Barrier $U(R_{C})$ relative to active energy scale as a function of activity and supersaturation. The distinct white contour lines distinguish values of $U(R_{C})/k_{B}T^{\rm act}$ separated by orders of magnitude.
As our deterministic force is a function of a single variable, $R$, we can define (within an integration constant) an effective conservative potential $U(R)$:
$U(R)=8\pi\upgamma_{\rm ost}\left[\frac{R^{2}}{2}-\frac{R^{3}}{3R_{C}}\right]$ (12)
with $F^{D}(R)=-\partial U/\partial R$. In defining this potential, we have treated $\upgamma_{\rm ost}$ as independent of $R$ (just as in Ref. [4]), an approximation that appears to be well justified while $R\sim R_{C}$ [33]. This potential may be rearranged into a form familiar to equilibrium intuition [33]:
$U(R)=\underbrace{4\pi R^{2}\upgamma_{\rm ost}}_{\rm{Interface}}-\underbrace{\frac{4}{3}\pi R^{3}\mathcal{V}}_{\rm{Bulk}},$ (13a)
where:
$\mathcal{V}=\frac{\mathcal{P}(\rho^{\rm sat})\Delta\rho}{\rho^{\rm sat}}-\frac{\Delta\Phi\Delta\rho}{\rho^{\rm sat}\Delta\mathcal{E}}.$ (13b)
At equilibrium, $\mathcal{V}$ simplifies to the expected $\mu(\rho^{\rm sat})\Delta\rho-\Delta f$ [33]. From Eq. (13), one can clearly appreciate that the barrier contains two contributions: a positive term which depends solely on interfacial properties scaling as $R^{2}$ and a negative term which depends solely on bulk properties scaling as $R^{3}$. Thus the familiar picture of homogeneous nucleation where interfacial effects compete with bulk effects in order to define a critical radius and nucleation barrier is recovered but now contains generalized volumetric driving and interfacial penalty terms.
Evaluating the potential at the critical radius allows us to compare the relative strengths of the nucleation barrier across activities and saturations, as shown in Fig. 3. Here we see a decrease in the nucleation barrier as $\nu^{\rm sat}$ increases, consistent with equilibrium intuition. We further observe that the nucleation barrier height is much higher for droplets in comparison to bubbles. The disparity in nucleation barrier between droplets and bubbles has an attractive physical interpretation: much like the capillary tension of fluids comprised of ABPs [26], $\upgamma_{\rm ost}$ is rooted in nematic flows. As we discuss in recent work [26], these nematic flows lead to particles accumulating at convex liquid surfaces and escaping from concave liquid surfaces. Then for small nuclei $R<R_{C}$, where curvature is high, droplets (concave liquid surfaces) are less favorable to grow than bubbles (convex liquid surfaces). This demonstrates that while Eqs. (11) satisfies the fluctuation dissipation theorem and allows for the definition of a conservative potential, the underlying physics of this conservative potential remain rooted in nonequilibrium effects.
Figure 4: Surface tensions as defined by $\upgamma_{\rm mech}$, $k\to 0$ limit of $\upgamma_{\rm cw}$, and $R\to\infty$ limit of $\upgamma_{\rm ost}$ for both bubbles and droplets as a function of run length.
The recovery of an effective detailed balance with respect to the size of a sub-critical nucleus directly implies that the probability of observing a nucleus of size $R<R_{C}$ is proportional to ${\text{exp}\left[-U(R)/k_{B}T^{\rm act}\right]}$ [4]. We note that the potential defined here relied on the assumption that the density profiles and quantities such as $\upgamma_{\rm ost}$ are constant with $R$ an approximation that clearly breaks down when $R<<R_{C}$. Even in equilibrium, the nucleation potential formulated using bulk thermodynamic quantities (e.g  the so-called “capillary approximation”) is often used for all $R<R_{C}$ despite likely being beyond the range of applicability of bulk thermodynamic quantities [12, 48, 49]. Nevertheless, use of this potential for all $R$ allows one to straightforwardly extend the arguments of Becker-Döring theory [9, 12, 15] to extract the following kinetic expression for the rate at which an active droplet nucleates per unit volume [33]:
$\mathcal{R}=\frac{\left(\rho^{\rm sat}\right)^{2}}{\rho^{\rm in}}U_{o}\sqrt{\frac{\pi\upgamma_{\rm ost}}{2k_{B}T^{\rm act}}}\text{exp}\left[-\frac{U(R_{C})}{k_{B}T^{\rm act}}\right].$ (14)
The characteristic timescale to observe a nucleation event is given by $\tau_{\rm nuc}=(V\mathcal{R})^{-1}$ where $V$ is the total volume of the system. Thus the work of Redner et al. [2], which inferred the existence of a nucleation barrier by observing clustering statistics consistent with Becker-Döring theory from particle simulation, is well-justified on theoretical grounds. We note that the rate [Eq. (14)] is applicable to the nucleation of active droplets but not to active bubbles – the physical interpretation of the “clusters” posited in the derivation of Eq. (14) becomes unclear [12] in this context.
II.5 Active Coarsening Dynamics
While Eqs. (6) and Eqs. (11) were motivated to describe the earliest stages of nucleation, i.e. the structure and dynamics of a single nucleus forming in a supersaturated solution, they may also describe late-stage coarsening dynamics long after all nucleation events have occurred. At this stage, the system will contain several nuclei of different sizes. The coarsening of these domains can proceed through a combination of several different mechanisms, including evaporation-condensation (Ostwald ripening) and coalescence. At sufficiently low $\nu^{\rm sat}$, the nuclei will be separated by large enough distances that collisions between them are exceedingly rare and coarsening will be dominated by Ostwald ripening. In such a scenario, Eqs. (11) can describe the dynamics of each nucleus such that they are only coupled via the density of the majority phase, $\rho^{\rm sat}$, which now evolves with time. When $\upgamma_{\rm ost}>0$, we expect the forward Ostwald ripening process to occur and larger nuclei of minority phase to grow at the expense of smaller nuclei until the system reaches a final steady state with a single macroscopic domain. Nuclei compete to absorb (or inject) particles from the majority phase and therefore $R_{C}$, the radius below which nuclei shrink and above which nuclei grow, is now a function of time.
With the above physical picture in mind, we can determine the time-dependence of the post-nucleation critical radius in a similar manner as the Lifshitz-Slyozov-Wagner (LSW) theory [11]. In the limit of a vanishing phase fraction of the minority phase, the nuclei can be approximated as independent with each satisfying the Langevin equation shown in Eq. (11). The Fokker-Planck equation consistent with Eq. (11) can be straightforwardly found and its long-time asymptotic solution (while discarding terms of higher order than $R^{-3}$) can be determined analogously to LSW theory, as detailed in Section 4 of the SI [33]. In doing so, we find that the critical radius evolves in time according to:
$R_{C}(t)=\left[R^{3}_{C}(t_{o})+\frac{6\rho^{\rm sat}}{\zeta\Delta\rho^{2}}\upgamma_{\rm ost}\mathcal{C}t\right]^{1/3},$ (15)
where $\mathcal{C}$ is a dimensionless constant and $t_{o}$ is an arbitrary time origin. We note that this long-time asymptotic solution also implies that the average radius is always within a dimensionless multiplicative constant of $R_{C}$ [33, 11, 14]. Equation (15) demonstrates that the $t^{1/3}$ growth law predicted by LSW theory should generically hold out of equilibrium while $\upgamma_{\rm ost}>0$. This $1/3$ power law has been measured experimentally in active systems [6, 7] with disparate microscopic details, suggesting that this growth law may in fact be robust. We note that the simulations of Redner et al. [8] measured the average number of particles per droplet (in two-dimensions) to grow slightly faster than $t^{1/2}$. This is consistent with a radial growth rate faster than $t^{1/4}$ and is perhaps consistent with our theoretical prediction.
We also find that Eqs. (6) may be used to treat coarsening problems with a finite phase fraction of minority phase, which is beyond the scope of LSW theory. To demonstrate this, we adopt the perspective and numerical scheme of Voorhees and Glicksman [31, 32], originally developed to study precipitation of solid phases in alloys. We now consider timescales at which the minority phase has nucleated and coarsened to the extent that the final phase fraction has been reached. All that remains to complete the phase separation process is for the disparate nuclei to coarsen to a single domain such that a single interface exists.
In the numerical scheme, the initial sizes and positions of nuclei in a fully periodic domain are prescribed. While the positions remain fixed, the sizes of the nuclei evolve with time via diffusion of particles through the majority phase. These effective diffusive interactions between nuclei, responsible for Ostwald ripening, can be calculated analogously to the electrostatic energy of a collection of point charges [31, 33]. This perspective, while initially developed for passive systems, only relies on thermodynamic assumptions by employing an equilibrium Gibbs-Thomson relation as a boundary condition between each nucleus and the majority phase. Our derived Eqs. (6) are the analogue of the Gibbs-Thomson relation for an active system, allowing us to extend the perspective of Ref. [31] to active matter.
For demonstration, we implement the active Voorhees-Glicksman theory to simulate the ripening of a periodic cell of active nuclei. Simulations of $250$ nuclei with random positions [50] and sizes were executed [33] with the nuclei separated by exceedingly large distances. It is therefore expected for the coarsening dynamics to be in agreement with LSW theory [11]. Noting again that the average radius is proportional to the critical radius, we expect the average radius across multiple activities to collapse to the same curve when scaled by the prefactor of the growth rate found in Eq. (15), ${K\equiv\left(\rho^{\rm sat}\upgamma_{\rm ost}d_{\rm hs}/\zeta U_{o}\Delta\rho^{2}\right)^{1/3}}$. The average nucleus size from these simulations scaled by $K$ are shown in Fig. 5. Our numerical simulations reveal that the average nucleus size is an excellent agreement with the predicted $\sim t^{1/3}$ scaling with curves for both bubbles and droplets collapsing for all activities. Plots of the distribution of nucleus sizes, which display evidence of evolving towards an LSW distribution [11], are provided in Section 4 of the SI [33].
Figure 5: Average radius growth as a function of time for nuclei initialized with an LSW distribution [11] and average radius $750d_{\rm hs}$. Radius growth curves across different activities, as well as between bubbles and droplets, collapse onto one another when scaled by prefactor ${K\equiv\left(\rho^{\rm sat}\upgamma_{\rm ost}d_{\rm hs}/\zeta U_{o}\Delta\rho^{2}\right)^{1/3}}$.
Although the Voorhees-Glicksman theory can be conveniently applied to late-stage active coarsening dynamics via Eqs. (6), it has significant limitations. In particular, the positions of the nuclei are fixed in time while in reality nuclei may diffuse, which may have significant impact on the inter-nuclei diffusion interactions and in extreme cases may lead to collision of nuclei (coalescence). More sophisticated coarsening theories [13] attempting to account for these effects, such as the work of Enomoto et al. [51] also only draw on thermodynamic arguments in the form of a Gibbs-Thomson relation. Using our work, such theories may then be extended to active matter so long as the non-equilibrium analogue to the Gibbs-Thomson relation employed in those works is established.
III Discussion and conclusions
We derive a general scheme for computing the properties of a critical nucleus arbitrarily far from equilibrium and apply this perspective to motility-induced phases comprised of active Brownian particles. From this theory we are able to derive generalized Gibbs-Thompson equations for active systems, a result which has been measured numerically but lacked theoretical justification [22]. Expanding on this theory, we use the framework proposed by Cates and Nardini [4] to find the Langevin dynamics of a nucleus. These dynamics allow us to define an effective nucleation barrier, revealing that the kinetic barrier to nucleating a droplet is substantially higher than that of a bubble. In addition, we find that the critical radius is an unstable fixed point for all activities where motility-induced phase separations occurs, demonstrating the nucleation dynamics of ABPs occur similarly to a passive system. By following the arguments first presented by Becker and Döring [9], we also determine the expected rate of nucleation for active droplets [2]. Our generalization of a Gibbs-Thompson relation to active systems allows a wide body of existing work studying the coarsening dynamics [11, 13] of passive systems to be applied to active matter. As an example, we demonstrated the applicability of Voorhees-Glicksman theory [31, 32] to active Brownian particles.
Our theory identifies physical implications of three independently defined surface tensions, shown in Fig. 4. The mechanical tension, $\upgamma_{\rm mech}$, which can be negative [21], controls the pressure difference between a nucleus and its surroundings only in the absence of non-local contributions to the dynamic stress. The capillary tension, $\upgamma_{\rm cw}$, establishes a surface area minimizing principle for an interface, which allows us to approximate fluctuating nuclei as spherically symmetric [25, 4, 26]. The Ostwald tension, $\upgamma_{\rm ost}$, plays a role in the finite-sized coexistence criteria in addition to controlling the dynamics of a nucleus [23, 4]. All three of these surface tensions are identical and positive for an equilibrium system but may take disparate, and even negative, values in active systems.
The controversy surrounding a negative $\upgamma_{\rm mech}$ has sparked a number of new definitions of surface tension in non-equilibrium systems. These methods have been rooted in a variety of perspectives, ranging from mechanical balances [52, 53] to dynamics [22, 25] to power functional theory [24]. Here, we argue that outside of equilibrium the surface tension must be defined contextually. While $\upgamma_{\rm mech}$ and $\upgamma_{\rm ost}$ were not found to enter the dynamics of capillary-waves [26], they are crucial for describing nucleation. Therefore one cannot haphazardly apply the definition of surface tension from one context to another; the physically relevant surface tension must be derived from first principles. While this work provides insight to the ongoing debates surrounding nonequilibrium surface tension, it may also serve as a starting point for other puzzles in active matter. This work could be extended to include multiple order parameters [44] in order to understand the nucleation of active crystals. Furthermore, applying the perspective of this work, in conjunction with the fluctuating hydrodynamics [26], may lead to identification of microscopic systems where $\upgamma_{\rm ost}$ changes sign and novel nucleation dynamics are uncovered [23].
Nucleation events are intrinsically challenging to observe and characterize in simulation and experiment. Indeed, measuring the nucleation rates of equilibrium systems as simple as hard spheres has been a longstanding challenge [54, 55, 56, 57, 58]. Theories for nucleation are therefore essential for both guiding computational and experimental investigations and making predictions for conditions in which measurements can be prohibitive. To this end, CNT has provided crucial intuition for nucleation studies. It is our hope that this work, which makes use of a number of important recent developments [23, 30, 4, 25, 26], can provide a similar road map for studies of active nucleation by extending CNT and canonical coarsening theories to systems far from equilibrium.
Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, under Award Number DE-SC0024900. L.L. was supported in part by the Department of Defense (DoD) through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program.

### v2

@@
 We briefly recapitulate the thermodynamic arguments of CNT applied to bubble/droplet nucleation, the focus of this work.
-The complete derivation can be found in Section 3 of the Supporting Information (SI)  [^1] See Supporting Information at [URL].
+The complete derivation can be found in Section 3 of the Supplementary Material (SM).
 We consider a macroscopic homogeneous single-component system prepared with a uniform number density $\rho^{\rm sat}$ within a fixed system volume $V$ (the total number of particles is thus also fixed at $N=V\rho^{\rm sat}$).
 We focus on saturation densities $\rho^{\rm sat}$ within the liquid-gas binodal of the phase diagram but far from the spinodal boundaries, limiting our analysis to be far from the critical point.
+This limitation is necessary in order to ensure that the pathway of the phase transition is unambiguously nucleation and growth, rather than spinodal decomposition.
 In this region of the phase diagram, a system prepared with a spatially uniform density of $\rho^{\rm sat}$ will remain linearly stable against density fluctuations until a critical nucleus with density $\rho^{\rm in}$ and volume $V^{\rm nuc}$ is formed.
@@
 The critical radius is then simply ${R_{C}=2\upgamma/(\mu^{\rm sat}\Delta\rho-\Delta f})$ where the difference between an arbitrary property, $a$, in the interior and exterior of the nucleus is defined as $\Delta a\equiv a^{\rm in}-a^{\rm sat}$.
-The free energy barrier associated with the critical nucleus can be expressed as the sum of the volumetric and interfacial contributions ${F^{\rm{\dagger}}=4\pi R_{C}^{2}\upgamma-\frac{4}{3}\pi R_{C}^{3}(\mu^{\rm sat}\Delta\rho-\Delta f)}$ or equivalently, ${F^{\rm{\dagger}}=\frac{4}{3}\pi\upgamma R_{C}^{2}}$ [33].
+The free energy barrier associated with the critical nucleus can be expressed as the sum of the volumetric and interfacial contributions ${F^{\rm{\dagger}}=4\pi R_{C}^{2}\upgamma-\frac{4}{3}\pi R_{C}^{3}(\mu^{\rm sat}\Delta\rho-\Delta f)}$ or equivalently, ${F^{\rm{\dagger}}=\frac{4}{3}\pi\upgamma R_{C}^{2}}$.
 The probability of observing such a critical nucleus is thus expected to be $P(R_{C})\sim\text{exp}\left[F^{{\dagger}}/k_{B}T\right]$.
@@
 (1) are the two equations that allow for the determination of the two unknowns: $\rho^{\rm in}$ and $R_{C}$.
-In the following section, we derive analogous equations capable of determining $\rho^{\rm in}$ and $R_{C}$ for systems arbitrarily far from equilibrium by adopting a mechanical perspective.
+In Sec.
+II.2, we derive analogous equations capable of determining $\rho^{\rm in}$ and $R_{C}$ for systems arbitrarily far from equilibrium by adopting a mechanical perspective.
 II.2 Mechanics of a Critical Nucleus
@@
 Let us consider the dynamical definition of a critical nucleus.
-When the critical nucleus is an unstable fixed point  [^2] Recently Tjhung et al.
-[23] demonstrated in some active systems the critical nucleus may be a stable fixed point, in such a case nuclei smaller (larger) than $R_{C}$ will grow (shrink)., nuclei smaller (larger) than $R_{C}$ will be associated with a particle flux normal to the interface such that the nuclei shrink (grow).
+When the critical nucleus is an unstable fixed point [33], nuclei smaller (larger) than $R_{C}$ will be associated with a particle flux normal to the interface such that the nuclei shrink (grow).
 $R_{C}$ thus represents a critical size at which the direction of this flux switches, so the system must have zero flux normal to the interface precisely when a nucleus has a radius of $R_{C}$.
-For a single realization of an active fluid, the stochastic noise plays a crucial role in the density flux, which we will discuss further in the next section.
+For a single realization of an active fluid, the stochastic noise plays a crucial role in the density flux, which we will discuss further in Sec.
+II.4.
 For the purposes of determining the properties of the critical nucleus, we can consider an ensemble of supersaturated states.
@@
 The critical nucleus will correspond to an additional state in which a finite-size minority phase with zero radial flux coexists with a parent phase.
-The condition for determining a stationary state of a density profile is given by the linear momentum balance (see Section 1 of the SI [33] and Ref.
+The condition for determining a stationary state of a density profile is given by the linear momentum balance (see Section 1 of the SM and Ref.
 [34]) which governs the dynamics of the density flux.
@@
 We thus focus on systems in which $\bm{\sigma}$ and $\mathbf{b}$ can be fully expressed as local density gradient expansions.
+Such a focus may become limiting for application to systems with truly long-ranged forces that prevent a local expansion of stresses/body forces.
 Up to a divergence-free term, we may absorb the body force into a stress-like contribution such that ${\bm{\nabla}\cdot\bm{\sigma}+\mathbf{b}=\bm{\nabla}\cdot\left[\bm{\sigma}+\mathbf{F}\right]=\bm{\nabla}\cdot\bm{\Sigma}}$, where $\mathbf{F}$ is the Coulomb integral of $\mathbf{b}$ and $\bm{\Sigma}\equiv\bm{\sigma}+\mathbf{F}$ is the dynamic stress tensor.
@@
 Under spherical symmetry, $\bm{\Sigma}$ has only two independent components and can be expressed as $\bm{\Sigma}=\Sigma_{rr}\mathbf{e}_{r}\mathbf{e}_{r}+\Sigma_{tt}\left(\mathbf{I}-\mathbf{e}_{r}\mathbf{e}_{r}\right)$ where $\mathbf{I}$ is the identity tensor.
-We can make analytical progress by expanding the dynamic stress tensor to second order in density gradients while leaving any potential non-local contributions symbolic:
-$\Sigma_{rr}=$ $-\mathcal{P}(\rho)+\kappa_{1}(\rho)\left(\frac{\partial}{\partial r}+\frac{2}{r}\right)\frac{\partial\rho}{\partial r}$
-$+\bigl{(}\kappa_{2}(\rho)+\kappa_{3}(\rho)\bigr{)}\left(\frac{\partial\rho}{\partial r}\right)^{2}+\kappa_{4}(\rho)\frac{\partial^{2}\rho}{\partial r^{2}}+\mathcal{K}_{r},$ (2a)
-$\Sigma_{tt}=-\mathcal{P}(\rho)+\kappa_{1}(\rho)\left(\frac{\partial}{\partial r}+\frac{2}{r}\right)\frac{\partial\rho}{\partial r}+\kappa_{3}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\mathcal{K}_{t},$ (2b)
-where $-\mathcal{P}(\rho)$ is the bulk isotropic contribution to the dynamic stress tensor, $\mathcal{K}_{r/t}$ are the non-local terms originating from the Coulomb integral of $\mathbf{b}$, and $\{\kappa_{i}\}$ are interfacial coefficients.
+We can make analytical progress by expanding the dynamic stress tensor to second order in density gradients and assuming spatial isotropy in the expansion coefficients:
+$\Sigma_{tt}=-\mathcal{P}(\rho)+\kappa_{1}(\rho)\left(\frac{\partial}{\partial r}+\frac{2}{r}\right)\frac{\partial\rho}{\partial r}+\kappa_{2}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\mathcal{K},$ (2b)
+where $-\mathcal{P}(\rho)$ is the bulk isotropic contribution to the dynamic stress tensor, $\mathcal{K}$ is the non-local term originating from the Coulomb integral of $\mathbf{b}$, and $\{\kappa_{i}\}$ are interfacial coefficients.
 These interfacial coefficients can be found by formally carrying out the expansion of $\bm{\sigma}$ and $\mathbf{b}$ beginning from the microscopic definition of stress and expanding the local terms that arise from the Coulomb integral of $\mathbf{b}$.
-The procedure for determining the precise forms of $\{\kappa_{i}\}$ and $\mathcal{K}_{r/t}$ from the expansion coefficients of $\bm{\sigma}$ and $\mathbf{b}$ is outlined in Section 1 of the SI [33].
+The procedure for determining the precise forms of $\{\kappa_{i}\}$ and $\mathcal{K}$ from the expansion coefficients of $\bm{\sigma}$ and $\mathbf{b}$ is outlined in Section 1.7 of the SM.
+In particular, a form for the nonlocal $\mathcal{K}$ term is found.
+We note that this form is free of any terms with explicit dependence on $1/r$.
 In a passive system $\mathcal{P}(\rho)$ is simply the bulk pressure and $\{\kappa_{i}\}$ are given by the Korteweg expansion [39] while non-local terms can be well-approximated by local terms for single-component systems with short-ranged interactions.
 More generally, the bulk pressure and interfacial coefficients can contain dynamical contributions otherwise absent in passive systems [26].
-Given the above physical picture, we now outline our process to find the zero radial flux states associated with a finite-sized minority phase [33].
+We also note that under the spherically symmetric configuration and stresses assumed here, setting the radial flux to zero also implies zero flux in the angular directions.
+In two dimensions, it is possible for a critical nucleus with radial symmetry to develop a finite angular flux if the underlying microscopic dynamics violate spatial parity.
+Describing the critical nucleus in chiral active matter [40, 41], would require a different expansion of the stress tensor such as those presented in Refs.
+[42, 43].
+Given the above physical picture, we now outline our process to find the zero radial flux states associated with a finite-sized minority phase.
 Our spherically symmetric geometry results in the linear momentum balance simplifying to:
@@
 (3) from $r=0$ to $r=\infty$ results in:
-$\Delta\mathcal{P}=\Delta\mathcal{K}_{r}+\frac{2\upgamma_{\rm mech}}{R_{C}}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right).$ (4a)
+$\Delta\mathcal{P}=\mathcal{K}\bigr{|}_{r=0}+\frac{2\upgamma_{\rm mech}}{R_{C}}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right).$ (4a)
 Here, $\upgamma_{\rm mech}$ is the mechanical surface tension as defined by Kirkwood and Buff [44]:
 $\upgamma_{\rm mech}$ $=\int_{0}^{\infty}\left[\Sigma_{tt}-\Sigma_{rr}\right]\text{d}r,$
-$=-\int_{0}^{\infty}\left[\kappa_{3}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\kappa_{4}(\rho)\frac{\partial^{2}\rho}{\partial r^{2}}+\mathcal{K}_{r}-\mathcal{K}_{t}\right]\text{d}r,$ (4b)
+$=-\int_{0}^{\infty}\left[\kappa_{3}(\rho)\left(\frac{\partial\rho}{\partial r}\right)^{2}+\kappa_{4}(\rho)\frac{\partial^{2}\rho}{\partial r^{2}}\right]\text{d}r,$ (4b)
 where the second equality utilizes our specific form of the dynamic stress [see Eqs.
@@
 Equation (4a) is a generalized Young-Laplace equation – with the equilibrium surface energy and pressure more generally replaced by the mechanical tension and dynamic pressure, respectively, and an additional contribution arising from the non-local terms.
+Equation (4) carries the familiar mechanical intuition of surface tension: it measures the excess stress parallel to the interface.
 This mechanical tension was determined to be strongly negative for phase-separated ABPs [21] and has been the subject of some controversy regarding its physical relevance.
@@
 [30] correctly identified that the mechanical surface tension of Kirkwood and Buff controls the Laplace pressure difference of droplets consistent with a fully local (dynamic) stress.
-More generally, $\Delta\mathcal{K}_{r}$ may also contribute to the Laplace pressure difference.
+More generally, the nonlocal term $\mathcal{K}$ may also contribute to the Laplace pressure difference.
+Note that Eq.
+(4a) ignores terms higher order in $1/R_{C}$.
+This approximation becomes increasingly accurate as the saturation is reduced [see Fig.
+2].
 While the generalized Young-Laplace equation provides a mechanical condition remarkably similar to its equilibrium counterpart, we still require a second condition to allow us to determine our two unknowns, $\rho^{\rm in}$ and $R_{C}$.
@@
 The search for chemical potential-like quantities for nonequilibrium systems has been the subject of much interest in the last decade in the context of determining macroscopic binodals [19, 45, 20, 34].
-These works have resulted in generalized Maxwell constructions [20, 35] that are derived from performing the appropriate integration of the mechanical criterion of coexistence [43, 35].
+These works have resulted in generalized Maxwell constructions [20, 34] that are derived from performing the appropriate integration of the mechanical criteria of coexistence [46, 34].
 We follow similar arguments used in constructing the nonequilibrium coexistence criteria of macroscopic binodals [46, 34].
@@
 We then introduce the generalized Maxwell construction pseudovariable [46, 20, 34] $\mathcal{E}$ defined by the differential equation ${(\kappa_{1}+\kappa_{4})\partial^{2}\mathcal{E}/\partial\rho^{2}=(2(\kappa_{2}+\kappa_{3})-\partial(\kappa_{1}+\kappa_{4})/\partial\rho)\partial\mathcal{E}/\partial\rho}$, which is found to be $\mathcal{E}\sim 1/\rho$ for a passive system.
-We multiply the indefinite integral of Eq.
-(3) by $d\mathcal{E}/dr$ and integrate from $r=0$ to $r=\infty$ to arrive at:
-$\int_{\mathcal{E}^{\rm in}}^{\mathcal{E}^{\rm sat}}\text{d}\mathcal{E}$ $\left[\mathcal{P}(\rho)-\mathcal{P}(\rho^{\rm sat})\right]=\frac{2\upgamma_{\rm ost}}{R_{C}}\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right),$ (5a)
+This pseudovariable allows us to eliminate nonlinear terms (just as it does in the case of macroscopic coexistence [46, 20, 34]) by multiplying the antiderivative of Eq.
+(3) by $d\mathcal{E}/dr$ and integrating from $r=0$ to $r=\infty$ to arrive at:
+$\int_{\mathcal{E}^{\rm in}}^{\mathcal{E}^{\rm sat}}\text{d}\mathcal{E}\left[\mathcal{P}(\rho)-\mathcal{P}(\rho^{\rm sat})\right]+$ $\int_{0}^{\infty}\left(\mathcal{E}^{\rm in}-\mathcal{E}\right)\frac{\partial\mathcal{K}}{\partial r}\text{d}r$
+$=$ $\frac{2\upgamma_{\rm ost}}{R_{C}}\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}+\mathcal{O}\left(\frac{1}{R_{C}^{2}}\right),$ (5a)
 where, analogously to Tjhung et al. [23], we have defined the “Ostwald” tension as
-$\upgamma_{\rm ost}=\frac{\Delta\rho}{\rho^{\rm sat}\Delta\mathcal{E}}$ $\Biggl{[}\int_{0}^{\infty}\Bigl{(}\kappa_{1}(\rho)\frac{\partial\mathcal{E}}{\partial\rho}\left(\frac{\partial\rho}{\partial r}\right)^{2}$
-$+\mathcal{G}$ $\left(\mathcal{E}-\mathcal{E}^{\rm in}\right)+\frac{R_{C}}{2}\frac{\partial\mathcal{K}_{r}}{\partial r}\left(\mathcal{E}^{\rm in}-\mathcal{E}\right)\Bigr{)}\text{d}r\Biggr{]}.$ (5b)
-Note that $\upgamma_{\rm ost}$ has been defined such that it has the same units as mechanical surface tension and becomes identical to $\upgamma_{\rm mech}$ in a passive system [33].
+$\upgamma_{\rm ost}=\frac{\Delta\rho}{\rho^{\rm sat}\Delta\mathcal{E}}\Biggl{[}\int_{0}^{\infty}\Bigl{(}\kappa_{1}(\rho)\frac{\partial\mathcal{E}}{\partial\rho}\left(\frac{\partial\rho}{\partial r}\right)^{2}+\mathcal{G}\left(\mathcal{E}-\mathcal{E}^{\rm in}\right)\Bigr{)}\text{d}r\Biggr{]}.$ (5b)
+Note that $\upgamma_{\rm ost}$ has been defined such that it has the same units as mechanical surface tension and becomes identical to $\upgamma_{\rm mech}$ in a passive system.
+$\upgamma_{\rm ost}$ is named for its importance in understanding the evaporation-condensation mechanism of active coarsening, which we will discuss in Sec.
+II.5.
+As $\rho^{\rm sat}$ approaches a binodal of the phase diagram, the right hand side of Eq.
+(5a) vanishes and we are left with one of the macroscopic coexistence criteria.
+Interestingly, a term depending on gradients (and therefore the interface) remains.
+At equilibrium, the details of the interface do not modify the coexistence criteria as interfacial effects are generally sub-extensive in the free energy.
+However, outside of equilibrium no such guarantee exists.
+This interfacial term vanishes as the derivative of $\mathcal{K}$ approaches zero, and one may refer to the expansion of $\mathcal{K}$ in the SM for a precise expression of this limit.
 The connection between Eqs.
@@
 It was subsequently shown that this potential satisfies a generalized Gibbs-Duhem relation [47], which, for the single-component system under consideration, is expressed as $\text{d}g=\mathcal{E}\text{d}\mathcal{P}$ where $g$ is the generalized chemical potential.
-The generalized Gibbs-Duhem relation allows [33] Eq.
+The generalized Gibbs-Duhem relation allows Eq.
 (5a) to be expressed as a difference of chemical potentials with the nonequilibrium critical nucleus criteria now taking the form:
-$\Delta g=\mathcal{E}^{\rm in}\Delta\mathcal{K}_{r}+\frac{2}{R_{C}}\left[\mathcal{E}^{\rm in}\upgamma_{\rm mech}-\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}\upgamma_{\rm ost}\right],$ (6a)
-$\Delta\mathcal{P}=\Delta\mathcal{K}_{r}+\frac{2\upgamma_{\rm mech}}{R_{C}}.$ (6b)
+$\Delta g=\mathcal{E}^{\rm in}\mathcal{K}\bigr{|}_{r=0}$ $+\int_{0}^{\infty}\left(2\mathcal{E}^{\rm in}-\mathcal{E}\right)\frac{\partial\mathcal{K}}{\partial r}\text{d}r$
+$+\frac{2}{R_{C}}\left[\mathcal{E}^{\rm in}\upgamma_{\rm mech}+\frac{\rho^{\rm sat}\Delta\mathcal{E}}{\Delta\rho}\upgamma_{\rm ost}\right],$ (6a)
+$\Delta\mathcal{P}=\mathcal{K}\bigr{|}_{r=0}+\frac{2\upgamma_{\rm mech}}{R_{C}}.$ (6b)
 Additionally, by defining a pseudopotential ${\Phi\equiv\int\mathcal{P}\text{d}\mathcal{E}}$, Eq.
@@
 One can use these relations to readily verify that $g^{\rm in}=g^{\rm sat}$ and that Eq.
-(5a) recovers the thermodynamically derived expression for $R_{C}$ [33].
+(5a) recovers the thermodynamically derived expression for $R_{C}$.
 Our mechanical criteria for the critical nucleus [Eqs.
@@
 In Ref.
-[23], similar finite-sized coexistence criteria were derived with two key differences.
-First, the Laplace pressure difference in Ref.
-[23] was found to be controlled by the tension controlling Ostwald ripening rather than $\upgamma_{\rm mech}$.
-Additionally, the chemical potentials (including non-local terms) inside and outside of the nucleus were found to be equal by Ref.
-[23].
-Fundamentally, this is because pressure and chemical potential have different physical interpretations in Ref.
+[23], finite-sized coexistence criteria were derived which appear to be distinct from those of the present work.
+Fundamentally, this is because pressure (referred to as “pseudopressure”), surface tension (referred to as “pseudotension”), and chemical potential have different physical interpretations in Ref.
 [23] than in this work.
-Here, pressure is the bulk isotropic component of the dynamic stress tensor and carries the familiar mechanical interpretation, while in Ref.
-[23] the pressure (also referred to as a “pseudopressure”) cannot be physically interpreted without identifying the pseudovariable.
-In Ref.
-[23] the chemical potential is defined as the scalar field controlling the dynamics and carries much of the familiar intuition of a chemical potential, while $g$ in this work cannot be physically interpreted without identifying $\mathcal{E}$.
-The mechanical approach presented above may be a more natural starting point when beginning with microscopic equations-of-motion and using an Irving-Kirkwood approach to derive the form of the dynamic stress and offers the advantage of clearly identifying the role of the mechanical tension of Kirkwood and Buff.
+Indeed, Ref.
+[23] explicitly cautioned against interpreting pseudopressure and pseudotension in the same manner as mechanical pressure and mechanical surface tension.
+In Section 1.6 of the SM we formally map AMB+ to our mechanical treatment of finite-size coexistence.
+Despite their ambiguous physical interpretation, we find that the finite-sized coexistence criteria presented in Ref.
+[23] are fully consistent with those found here.
+The unambiguous physical interpretation of the mechanical quantities used in our work represents an advantage of a mechanical perspective which also has a clear connection to microscopics through the Irving-Kirkwood-Noll procedure [35, 36, 37].
 Given a saturation density, one may use Eqs.
@@
 Therefore, solution of Eqs.
-(6) requires self-consistently solving for the density profile using a procedure detailed in Section 1 of the SI [33].
-In the following section, we apply our mechanical theory to a canonical model system for active matter, active Brownian particles.
+(6) requires self-consistently solving for the density profile using a procedure detailed in Section 1 of the SM.
+In Sec.
+II.3, we apply our mechanical theory to a canonical model system for active matter, active Brownian particles.
 II.3 Nucleation of Motility-Induced Phases
@@
 Here, the volume fraction is defined as the number density multiplied by the effective volume of a particle $\rho\pi d_{\rm hs}^{3}/6$.
-The form of $\mathcal{P}(\rho)$ and $\{\kappa_{i}(\rho)\}$ for ABPs are given by [35, 26]:
-$\mathcal{P}(\rho)$ $=p_{C}(\rho)+\frac{\rho\zeta U_{o}\ell_{o}\overline{U}(\rho)}{d(d-1)},$ (8a)
-$\kappa_{1}(\rho)$ $=\frac{3\ell_{o}^{2}}{2d(d-1)(d+2)}\left(\overline{U}(\rho)\right)^{2}\frac{\partial p_{C}}{\partial\rho},$ (8b)
-$\kappa_{3}(\rho)$ $=\frac{3\ell_{o}^{2}}{2d(d-1)(d+2)}\overline{U}(\rho)\frac{\partial}{\partial\rho}\left[\overline{U}(\rho)\frac{\partial p_{C}}{\partial\rho}\right],$ (8c)
+In three dimensions, the form of $\mathcal{P}(\rho)$ and $\{\kappa_{i}(\rho)\}$ for ABPs are given by [34, 26]:
+$\mathcal{P}(\rho)$ $=p_{C}(\rho)+\frac{\rho\zeta U_{o}\ell_{o}\overline{U}(\rho)}{6},$ (8a)
+$\kappa_{1}(\rho)$ $=\frac{\ell_{o}^{2}}{20}\left(\overline{U}(\rho)\right)^{2}\frac{\partial p_{C}}{\partial\rho},$ (8b)
+$\kappa_{3}(\rho)$ $=\frac{\ell_{o}^{2}}{20}\overline{U}(\rho)\frac{\partial}{\partial\rho}\left[\overline{U}(\rho)\frac{\partial p_{C}}{\partial\rho}\right],$ (8c)
 $\kappa_{2}(\rho)$ $=\kappa_{4}(\rho)=0,$ (8d)
-$\mathcal{K}_{r}$ $=\mathcal{K}_{t}=0,$ (8e)
-where $p_{C}(\rho)$ is the conservative pressure arising from particle interactions, $\overline{U}(\rho)$ is the dimensionless effective active speed, and $d(>1)$ is the system dimensionality.
+$\mathcal{K}$ $=0,$ (8e)
+where $p_{C}(\rho)$ is the conservative pressure arising from particle interactions and $\overline{U}(\rho)$ is the dimensionless effective active speed.
 In Eqs.
-(8) we have ignored the gradient contributions of the interaction potential to $\bm{\Sigma}$, which is justified when $\ell_{o}/d_{\rm hs}>>1$ for purely repulsive interactions [30, 35, 44].
+(8) we have ignored the gradient contributions of the interaction potential to $\bm{\Sigma}$, which is justified when $\ell_{o}/d_{\rm hs}\gg 1$ for purely repulsive interactions [30, 34, 47].
 Given our expressions for $\{\kappa_{i}(\rho)\}$, one can solve for the generalized Maxwell construction pseudovariable $\mathcal{E}\sim p_{C}(\rho)$ [34].
-We also note that because $\mathcal{K}_{r}=0$, Eq.
+We also note that because $\mathcal{K}=0$, Eq.
 (4a) reduces to the familiar Laplace pressure difference (i.e.
 no non-local corrections) and the negative sign of $\upgamma_{\rm mech}$ [21] implies that the pressure inside of a droplet is lower than the pressure outside of the droplet [30].
-The phase diagram, dependent on $\ell_{o}/d_{\rm hs}$ and volume fraction $\phi\equiv\pi d_{\rm hs}^{3}/6$, has been well established computationally [5] and theoretically [35] with the latter predicting a critical point at $\phi\approx 0.49$ and $\ell_{o}\approx 16.2d_{\rm hs}$
-Figure 2: Critical radius and relative change in nucleus density (${\nu^{\rm in}\equiv|\rho^{\rm in}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$) as a function of supersaturation (${\nu^{\rm sat}\equiv|\rho^{\rm sat}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$).
+The phase diagram, dependent on $\ell_{o}/d_{\rm hs}$ and volume fraction $\phi\equiv\rho\pi d_{\rm hs}^{3}/6$, has been well established computationally [5] and theoretically [34] with the latter predicting a critical point at $\phi\approx 0.49$ and $\ell_{o}\approx 16.2d_{\rm hs}$
+Figure 2: Critical radius and relative change in nucleus density (${\nu^{\rm in}\equiv|\rho^{\rm in}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$) as a function of supersaturation (${\nu^{\rm sat}\equiv|\rho^{\rm sat}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$), as calculated from Eq.
+(6) using the numerical procedure outlined in Section 1.3 of the SM.
+Each data point requires self-consistent solution of the corresponding full spatial profile, e.g.
+those shown Fig.
+1.
+Linear dependence of $\nu^{\rm in}$ on $\nu^{\rm sat}$ and inverse dependence of $R_{C}/d_{\rm hs}$ on $\nu^{\rm sat}$ implies a Gibbs-Thomson relation with $\nu^{\rm{in}}\propto 1/R$
 With the dynamic stress of ABPs in hand, we are able to self-consistently solve our generalized critical nucleus conditions [Eqs.
@@
 This difference between bubbles and droplets widens as activity is increased and the phases become increasingly distinct.
-To contextualize this order of magnitude difference, we calculate the critical nucleus size and density for an equilibrium van der Waals fluid [33].
+To contextualize this order of magnitude difference, we calculate the critical nucleus size and density for an equilibrium van der Waals fluid.
 While the size of the critical droplets are indeed larger than the size of critical bubbles in passive fluids, they remain within an order of magnitude of each other at similar saturation conditions even far from the critical point.
@@
 3.
+For a more physical understanding of the discrepancy in critical nucleus size between bubbles and droplets, we draw on the kinetic arguments independently proposed by Redner et al.
+[2] and Lee [22].
+Establishing a flux-free critical nucleus with a density (and therefore interaction pressure) distinct from its surroundings requires the generation of an interfacial region of particles with a polarization pointing into the dense phase.
+Crucially, the rate at which particles with this orientation arrive to the interface must equal the rate at which they reorient and escape in order to satisfy the flux-free condition.
+The arrival rate is controlled by gas phase active speed and density while the escape rate is controlled by the minimum angle a particle must reorient in order to escape, the so-called “horizon angle [2].” For the same $R_{C}$, the rate of escape will depend greatly on whether the nucleus is the liquid phase (small horizon angle) or the gas phase (large horizon angle).
+As a result, for a similar supersaturation, the droplet must have a larger critical radius in order to satisfy the flux-free condition.
 II.4 Langevin Dynamics of a Nucleus
@@
 $\mathbf{J}$ $=\frac{1}{\zeta}\bm{\nabla}\cdot\bm{\Sigma}+\bm{\eta}^{\rm act},$ (9b)
-where $\rho$ now is a stochastic density field (while in the previous sections it was an ensemble average), $\bm{\Sigma}$ is defined identically as the dynamic stress tensor of the previous section and $\bm{\eta^{\rm act}}$ is an athermal stochastic contribution to the flux with zero mean and a variance of:
+where $\rho$ now is a stochastic density field (while in the previous sections it was an ensemble average), $\bm{\Sigma}$ is defined identically as the dynamic stress tensor of Sec.
+II.3 and $\bm{\eta^{\rm act}}$ is an athermal stochastic contribution to the flux with zero mean and a variance of:
 $\langle\bm{\eta}^{\rm act}(\mathbf{r},t)\bm{\eta}^{\rm act}(\mathbf{r}^{\prime},t^{\prime})\rangle=$ $2\frac{k_{B}T^{\rm act}}{\zeta}\left(\rho\mathbf{I}-\frac{d}{d-1}\mathbf{Q}^{\prime}\right)$
@@
 The ratio of the capillary tension to this energy scale – the active interfacial stiffness – controls the amplitude of long-wavelength interfacial fluctuations.
-A fluctuating nucleus can be well-approximated as a sphere when the average radius $R$ satisfies $(R^{2}\upgamma_{\rm cw}/k_{B}T^{\rm act})^{1/2}>>1$.
+A fluctuating nucleus can be well-approximated as a sphere when the average radius $R$ satisfies $(R^{2}\upgamma_{\rm cw}/k_{B}T^{\rm act})^{1/2}\gg 1$.
 Under this condition, the stochastic density field dynamics can be connected to the radial growth dynamics of the nucleus through the following radially symmetric ansatz [49, 4]:
@@
 In the case of a negative $\upgamma_{\rm ost}$, $R_{C}$ becomes a stable fixed point.
+In such a scenario, one may expect phenomonology ranging from microphase separation to a “bubbly liquid” phase coexisting with a gas phase [23].
 However, for ABPs we do not find a negative $\upgamma_{\rm ost}$ [see Fig.
-4] – implying that nucleation occurs with the same qualitative dynamics as an equilibrium system as observed in simulation [2].
+4] – consistent with existing simulation work on three-dimensional ABPs [51, 5, 52, 26].
+In two dimensions, some simulations of ABPs have reported phenomenology consistent with bubbly liquids [51, 53].
+Whether these bubbly phases are a result of a negative $\upgamma_{\rm ost}$ or other effects [54] remains an open question.
 Figure 3: Barrier $U(R_{C})$ relative to active energy scale as a function of activity and supersaturation.
@@
 In defining this potential, we have treated $\upgamma_{\rm ost}$ as independent of $R$ (just as in Ref.
-[4]), an approximation that appears to be well justified while $R\sim R_{C}$ [33].
-This potential may be rearranged into a form familiar to equilibrium intuition [33]:
+[4]), an approximation that appears to be well justified while $R\sim R_{C}$.
+This potential may be rearranged into a form familiar to equilibrium intuition:
 $U(R)=\underbrace{4\pi R^{2}\upgamma_{\rm ost}}_{\rm{Interface}}-\underbrace{\frac{4}{3}\pi R^{3}\mathcal{V}}_{\rm{Bulk}},$ (13a)
@@
 $\mathcal{V}=\frac{\mathcal{P}(\rho^{\rm sat})\Delta\rho}{\rho^{\rm sat}}-\frac{\Delta\Phi\Delta\rho}{\rho^{\rm sat}\Delta\mathcal{E}}.$ (13b)
-At equilibrium, $\mathcal{V}$ simplifies to the expected $\mu(\rho^{\rm sat})\Delta\rho-\Delta f$ [33].
+At equilibrium, $\mathcal{V}$ simplifies to the expected $\mu(\rho^{\rm sat})\Delta\rho-\Delta f$.
 From Eq.
@@
 We note that the potential defined here relied on the assumption that the density profiles and quantities such as $\upgamma_{\rm ost}$ are constant with $R$ an approximation that clearly breaks down when $R<<R_{C}$.
-Even in equilibrium, the nucleation potential formulated using bulk thermodynamic quantities (e.g  the so-called “capillary approximation”) is often used for all $R<R_{C}$ despite likely being beyond the range of applicability of bulk thermodynamic quantities [12, 48, 49].
-Nevertheless, use of this potential for all $R$ allows one to straightforwardly extend the arguments of Becker-Döring theory [9, 12, 15] to extract the following kinetic expression for the rate at which an active droplet nucleates per unit volume [33]:
+Even in equilibrium, the nucleation potential formulated using bulk thermodynamic quantities (e.g  the so-called ”capillary approximation”) is often used for all $R<R_{C}$ despite likely being beyond the range of applicability of bulk thermodynamic quantities [12, 55, 56].
+Nevertheless, use of this potential for all $R$ allows one to straightforwardly extend the arguments of Becker-Döring theory [9, 12, 15] to extract the following kinetic expression for the rate at which an active droplet nucleates per unit volume, as outlined in Appendix A:
 $\mathcal{R}=\frac{\left(\rho^{\rm sat}\right)^{2}}{\rho^{\rm in}}U_{o}\sqrt{\frac{\pi\upgamma_{\rm ost}}{2k_{B}T^{\rm act}}}\text{exp}\left[-\frac{U(R_{C})}{k_{B}T^{\rm act}}\right].$ (14)
@@
 Thus the work of Redner et al.
-[2], which inferred the existence of a nucleation barrier by observing clustering statistics consistent with Becker-Döring theory from particle simulation, is well-justified on theoretical grounds.
+[2], which inferred the existence of a nucleation barrier by observing clustering statistics consistent with Becker-Döring theory from particle simulation, is well-justified on theoretical grounds, although our calculated barrier is for three dimensions while the simulations of Ref.
+[2] are in two.
 We note that the rate [Eq.
@@
 (11) can describe the dynamics of each nucleus such that they are only coupled via the density of the majority phase, $\rho^{\rm sat}$, which now evolves with time.
-When $\upgamma_{\rm ost}>0$, we expect the forward Ostwald ripening process to occur and larger nuclei of minority phase to grow at the expense of smaller nuclei until the system reaches a final steady state with a single macroscopic domain.
+The finite-size coexistence criteria presented in this work imply minor differences in density of the majority phase just outside nuclei of different sizes, inversely proportional to their radii.
+The majority phase therefore has an inhomogeneous density, driving transport of mass away from nuclei with enriched outside densities towards nuclei with depleted outside densities.
+For passive systems as well as ABPs [see Fig.
+1], smaller droplets have enriched outside densities while smaller bubbles have depleted outside densities.
+Noting that transport of mass away from a nucleus would cause a droplet to shrink and a bubble to grow, this implies that large nuclei of minority phase will grow at the expense of smaller nuclei until the system reaches a final steady state with a single macroscopic domain.
 Nuclei compete to absorb (or inject) particles from the majority phase and therefore $R_{C}$, the radius below which nuclei shrink and above which nuclei grow, is now a function of time.
+While we do not have profiles for a particular system in which $\upgamma_{\rm ost}<0$, we suspect they would reveal smaller droplets having depleted outside densities and smaller bubbles having enriched outside densities.
+This would result in smaller nuclei growing at the expense of larger ones, consistent with stable radial dynamics of Eq.
+(11).
 With the above physical picture in mind, we can determine the time-dependence of the post-nucleation critical radius in a similar manner as the Lifshitz-Slyozov-Wagner (LSW) theory [11].
@@
 The Fokker-Planck equation consistent with Eq.
-(11) can be straightforwardly found and its long-time asymptotic solution (while discarding terms of higher order than $R^{-3}$) can be determined analogously to LSW theory, as detailed in Section 4 of the SI [33].
+(11) can be straightforwardly found and its long-time asymptotic solution (while discarding terms of higher order than $R^{-3}$) can be determined analogously to LSW theory, as detailed in Appendix B.
 In doing so, we find that the critical radius evolves in time according to:
@@
 where $\mathcal{C}$ is a dimensionless constant and $t_{o}$ is an arbitrary time origin.
-We note that this long-time asymptotic solution also implies that the average radius is always within a dimensionless multiplicative constant of $R_{C}$ [33, 11, 14].
+We note that this long-time asymptotic solution also implies that the average radius is always within a dimensionless multiplicative constant of $R_{C}$ [11, 14].
 Equation (15) demonstrates that the $t^{1/3}$ growth law predicted by LSW theory should generically hold out of equilibrium while $\upgamma_{\rm ost}>0$.
@@
 We note that the simulations of Redner et al.
-[8] measured the average number of particles per droplet (in two-dimensions) to grow slightly faster than $t^{1/2}$.
+[8] measured the average number of particles per droplet (in two-dimensions) to grow slightly faster than $t^{1/2}$ at late times.
 This is consistent with a radial growth rate faster than $t^{1/4}$ and is perhaps consistent with our theoretical prediction.
+Curiously, Ref.
+[8] also measured a much faster coarsening rate at early times where nuclei are spaced closely.
+While in this work we do not consider coarsening driven by nucleus collision, it is well-established for passive systems to also result in a $t^{1/3}$ growth law [57].
+A transition from fast coarsening when nuclei are closely spaced to slow coarsening when nuclei are distant is eerily reminiscent of coarsening in systems with hydrodynamic interactions, in which collisions of nuclei (and subsequent relaxation to spherical geometry) induce flows that cause further collisions [58, 59].
+While standard ABP models do not consider hydrodynamic interactions, a detailed accounting of the polarization field at the early stages of coarsening may reveal non-diffusive contributions.
+Furthermore, if $\ell_{o}$ greatly exceeds the spacing between nuclei, particles can be ballistically exchanged between nuclei, an effect that is also not accounted for in our perspective.
+A systematic treatment of early-stage active coarsening would help bridge the gap in our understanding between nucleation and late-stage coarsening and is the subject of future work.
 We also find that Eqs.
@@
 While the positions remain fixed, the sizes of the nuclei evolve with time via diffusion of particles through the majority phase.
-These effective diffusive interactions between nuclei, responsible for Ostwald ripening, can be calculated analogously to the electrostatic energy of a collection of point charges [31, 33].
-This perspective, while initially developed for passive systems, only relies on thermodynamic assumptions by employing an equilibrium Gibbs-Thomson relation as a boundary condition between each nucleus and the majority phase.
+Additionally, the majority phase will relax much faster than the individual nuclei grow or shrink.
+We therefore consider the diffusive dynamics of only the majority phase (i.e., excluding the regions occupied by the nuclei) with the nuclei simply acting as particle generation terms (sources or sinks of particles).
+Considering only the majority phase dynamics implies that large density-gradient interfaces are ignored; we therefore discard gradient terms from $\bm{\Sigma}$ and arrive at the quasi-stationary expression for the majority phase’s density field (see Appendix C for further details):
+$0=\bm{\nabla}^{2}\mathcal{P}(\rho)+\zeta\sum_{i=1}^{N_{\rm nuc}}S_{i}(\mathbf{r}),$ (16)
+where $N_{\rm nuc}$ is the number of nuclei and $S_{i}(\mathbf{r})$ is a function quantifying its injection/absorption of particles into the majority phase.
+Using Eq.
+(16) to solve for the pressure field is analogous to solving for the electrostatic potential resulting from $N_{\rm nuc}$ charge densities.
+Approximating each nucleus as spherically symmetric allows one to treat the “charge densities” as “point charges”, i.e.
+$S_{i}(\mathbf{r})\sim\delta(\mathbf{r}-\mathbf{r}_{i})$.
+Then the effective diffusive interactions between nuclei, responsible for Ostwald ripening, can be calculated analogously to the electrostatic energy of a collection of point charges [31].
+This perspective, while initially developed for passive systems, only relies on thermodynamic assumptions by employing an equilibrium Gibbs-Thomson relation as a boundary condition between each nucleus and the majority phase, allowing for the determination of the magnitude of each $S_{i}$.
 Our derived Eqs.
@@
 For demonstration, we implement the active Voorhees-Glicksman theory to simulate the ripening of a periodic cell of active nuclei.
-Simulations of $250$ nuclei with random positions [50] and sizes were executed [33] with the nuclei separated by exceedingly large distances.
+Simulations of $250$ nuclei with random positions [60] and sizes were executed with the nuclei separated by exceedingly large distances.
 It is therefore expected for the coarsening dynamics to be in agreement with LSW theory [11].
@@
 Our numerical simulations reveal that the average nucleus size is an excellent agreement with the predicted $\sim t^{1/3}$ scaling with curves for both bubbles and droplets collapsing for all activities.
-Plots of the distribution of nucleus sizes, which display evidence of evolving towards an LSW distribution [11], are provided in Section 4 of the SI [33].
-Figure 5: Average radius growth as a function of time for nuclei initialized with an LSW distribution [11] and average radius $750d_{\rm hs}$.
+Plots of the distribution of nucleus sizes, which display evidence of evolving towards an LSW distribution [11], are provided in Section 4 of the SM.
+Figure 5: Average radius growth as a function of time for nuclei initialized with an LSW distribution and average radius $750d_{\rm hs}$.
 Radius growth curves across different activities, as well as between bubbles and droplets, collapse onto one another when scaled by prefactor ${K\equiv\left(\rho^{\rm sat}\upgamma_{\rm ost}d_{\rm hs}/\zeta U_{o}\Delta\rho^{2}\right)^{1/3}}$.
@@
 [61] also only draw on thermodynamic arguments in the form of a Gibbs-Thomson relation.
-Using our work, such theories may then be extended to active matter so long as the non-equilibrium analogue to the Gibbs-Thomson relation employed in those works is established.
+Our work establishes a Gibbs-Thomson relation for single-component phase separation purely from mechanics, without invoking equilibrium assumptions.
+Then established perspectives on classical coarsening may be extended to nonequilibrium systems as long as a nonequilibrium Gibbs-Thomson relation, such as that found in this work, is used.
 III Discussion and conclusions
-We derive a general scheme for computing the properties of a critical nucleus arbitrarily far from equilibrium and apply this perspective to motility-induced phases comprised of active Brownian particles.
+We derive a mechanical scheme for computing the properties of a critical nucleus arbitrarily far from equilibrium and apply this perspective to motility-induced phases comprised of active Brownian particles.
 From this theory we are able to derive generalized Gibbs-Thompson equations for active systems, a result which has been measured numerically but lacked theoretical justification [22].
@@
 By following the arguments first presented by Becker and Döring [9], we also determine the expected rate of nucleation for active droplets [2].
-Our generalization of a Gibbs-Thompson relation to active systems allows a wide body of existing work studying the coarsening dynamics [11, 13] of passive systems to be applied to active matter.
+Our generalization of a Gibbs-Thompson relation to mechanical nonequilibrium systems allows a wide body of existing work studying the coarsening dynamics [11, 13] of passive systems to be applied to active matter.
 As an example, we demonstrated the applicability of Voorhees-Glicksman theory [31, 32] to active Brownian particles.
@@
 All three of these surface tensions are identical and positive for an equilibrium system but may take disparate, and even negative, values in active systems.
-The controversy surrounding a negative $\upgamma_{\rm mech}$ has sparked a number of new definitions of surface tension in non-equilibrium systems.
+Our results point to several new avenues for future work.
+One interesting direction would be to find rigorous connections between $\upgamma_{\rm ost}$ and more palpable kinetic arguments such as those proposed by Redner et al.
+[8, 2] and Lee [22], potentially aided by measurements from particle based simulations.
+Another would be to thoroughly investigate the physical implications of the nonlocal stress term $\mathcal{K}$ identified here, which may lead to macroscopic coexistence criteria explicitly dependent on interfacial properties.
+The controversy surrounding a negative $\upgamma_{\rm mech}$ has sparked a number of new definitions of surface tension in nonequilibrium systems.
 These methods have been rooted in a variety of perspectives, ranging from mechanical balances [62, 63] to dynamics [22, 25] to power functional theory [24].
@@
 It is our hope that this work, which makes use of a number of important recent developments [23, 30, 4, 25, 26], can provide a similar road map for studies of active nucleation by extending CNT and canonical coarsening theories to systems far from equilibrium.
-Acknowledgements.
+IV Supplemental Material
+See Supplementary Material for detailed derivation of coexistence criteria, nucleation dynamics, and coarsening behavior.
+V Acknowledgments
 This material is based upon work supported by the U.S.
@@
 was supported in part by the Department of Defense (DoD) through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program.
+VI Author Declarations
+The authors have no financial conflicts of interests to report.
+Appendix A Becker-Döring Nucleation Rate
+The existence of an effective potential and the establishment of nuclei obeying an effective detailed balance allows for one to extend equilibrium Becker-Dp̈ring theory [9, 12, 15] to active particles in order to predict the rate of droplet nucleation.
+The nucleation event is characterized by the formation of a cluster with a critical number of particles, so we would like to determine the rate at which such critical clusters form.
+We will denote the number of clusters with $\ell$ particles per unit volume as $c_{\ell}$.
+Additionally, we will assume that clusters are formed and destroyed by monomer attachment and detachment.
+In other words, a size $(\ell+1)$ cluster is formed by the merging of a size $\ell$ cluster and a size $1$ cluster.
+The opposite process would be a size $\ell$ cluster being destroyed by splitting into a size $\ell-1$ cluster and a size $1$ cluster.
+The flux, $J_{\ell}$, in cluster-size space of $\ell$ clusters is then given by the difference between the rate at which droplets containing $\ell$ particles are formed and the rate at which they are destroyed:
+$J_{\ell}=c_{\ell-1}S_{\ell-1}\beta_{\ell-1}-c_{\ell}S_{\ell}\alpha_{\ell};\ell\geq 2.$ (17)
+Here $S_{\ell}$ is the surface area of an $\ell$ cluster, $\beta_{\ell}$ is the flux of single particles arriving on the surface of the cluster, and $\alpha_{\ell}$ is the flux of single particles leaving the surface of the cluster.
+The evolution of $c_{\ell}$ is then given by:
+$\frac{\partial c_{\ell}}{\partial t}=J_{\ell}-J_{\ell+1}.$ (18)
+We now need to distinguish between two useful limits.
+One limit, which we call steady state requires that all $\partial_{t}c_{\ell}=0$ but the individual $J_{\ell}$ may remain finite.
+The other limit, which we call quasi-equilibrium requires that all $\partial_{t}c_{\ell}=0$ but additionally requires that all individual $J_{\ell}=0$.
+The effective detailed balance established for nucleation dynamics implies that the dynamics which govern quasi-equilibrium must also lead to quasi-equilibrium, as Eq.
+(11a) obeys an effective time-reversal symmetry.
+This implies that the values of $\beta_{\ell}$ and $\alpha_{\ell}$ which are valid at quasi-equilibrium must also be valid in the steady-state.
+Cates and Nardini [4] established that, to linear approximation, the principle of detailed balance is recovered by nucleating clusters in active system even if no such principle exists microscopically.
+Then $\alpha_{\ell}$ is:
+$\alpha_{\ell}=\frac{c_{\ell-1}^{\rm eq}S_{\ell-1}\beta_{\ell-1}}{c_{\ell}^{\rm eq}S_{\ell}},$ (19)
+and the flux $J_{\ell}$ at steady-state may be expressed as:
+$J_{\ell}=c_{\ell-1}S_{\ell-1}\beta_{\ell-1}-c_{\ell}S_{\ell}\frac{c_{\ell-1}^{\rm eq}S_{\ell-1}\beta_{\ell-1}}{c_{\ell}^{\rm eq}S_{\ell}}.$ (20)
+This flux may be rewritten as:
+$J_{\ell}=\beta_{\ell-1}S_{\ell-1}c_{\ell-1}^{\rm eq}\left[\frac{c_{\ell-1}}{c_{\ell-1}^{\rm eq}}-\frac{c_{\ell}}{c_{\ell}^{\rm eq}}\right].$ (21)
+Note that in the steady state all $J_{\ell}$ with $\ell\geq 2$ are equal to $J_{\ell+1}$. Then we can label $J_{\ell}=\mathcal{R}$ and write:
+$\frac{\mathcal{R}}{\beta_{\ell-1}S_{\ell-1}c^{\rm eq}_{\ell-1}}=\left[\frac{c_{\ell-1}}{c_{\ell-1}^{\rm eq}}-\frac{c_{\ell}}{c_{\ell}^{\rm eq}}\right].$ (22)
+We now sum both sides of the equation from $\ell=2$ to $\ell=\Lambda+1$, where $\Lambda$ is some very large number (larger than the critical nucleus size).
+After summing, the right hand side of Eq.
+(22) will vanish except for the first term where $\ell=2$ and the last term where $\ell=\Lambda+1$:
+$\mathcal{R}=\left[\frac{c_{1}}{c_{1}^{\rm eq}}-\frac{c_{\Lambda+1}}{c_{\Lambda+1}^{\rm eq}}\right]\left[\sum_{\ell=1}^{\Lambda}\frac{1}{\beta_{\ell}S_{\ell}c_{\ell}^{\rm eq}}\right]^{-1}.$ (23)
+We now argue that since the bulk metastable phase (a dilute gas) exists overwhelmingly as single particles, so the quasi-equilibrium and steady-state concentration of single particles are indistinuishable and $c_{1}/c_{1}^{\rm eq}\approx 1$.
+Additionally, $\Lambda$ is larger than the size of the critical nucleus, so $c_{\Lambda+1}$ should be zero in the metastable steady state but finite at quasi-equilibrium.
+Then Eq.
+(23) reduces to:
+$\mathcal{R}=\left[\sum_{\ell=1}^{\Lambda}\frac{1}{\beta_{\ell}S_{\ell}c_{\ell}^{\rm eq}}\right]^{-1}.$ (24)
+Using the effective nucleation potential [4], we know the relative probability of observing a nucleus with radius $R$, and (given a spherical geometry), we therefore we have an expression for $c_{\ell}^{\rm eq}$ as:
+$c_{\ell}^{\rm eq}=\rho^{\rm sat}\text{exp}\left[\left(k_{B}T^{\rm act}\right)^{-1}U(R_{\ell})\right],$ (25)
+where $U(R_{\ell})$ is the effective potential evaluated at a radius corresponding to a sphere of $\ell$ particles.
+Switching from a summation to an integral, we can then approximate the nucleation rate as:
+$\mathcal{R}=\rho^{\rm sat}\left[\int_{1}^{\Lambda}\text{d}\ell\frac{1}{\beta_{\rm\ell}S_{\ell}}\text{exp}\left[\frac{U(R_{\ell})}{k_{B}T^{\rm act}}\right]\right]^{-1}.$ (26)
+We now would like to make the connection between radius and cluster size more explicit:
+$\ell=\frac{4}{3}\pi R^{3}\rho^{\rm in}.$ (27)
+Alternatively, one may rearrange Eq. (27) and obtain an expression for the radius in terms of size:
+$R=\left[\frac{3\ell}{4\pi\rho^{\rm in}}\right]^{1/3}.$ (28)
+We now consider the derived form of the nucleation barrier, Eq. (13), and express it in terms of cluster size:
+$U(\ell)=4\pi\upgamma_{\rm ost}\left[\frac{3\ell}{4\pi\rho^{\rm in}}\right]^{2/3}-\frac{\ell\mathcal{V}}{\rho^{\rm in}}.$ (29)
+From this form the critical cluster size $\ell^{*}$ can be solved for by extremizing $U(\ell)$, resulting in:
+$\ell^{*}=\frac{32}{3}\pi\rho^{\rm in}\left(\frac{\upgamma_{\rm ost}}{\mathcal{V}}\right)^{3}.$ (30)
+We now Taylor expand the the barrier about this maximum to second order,
+$U(\ell)=U(\ell^{*})-\frac{1}{2}\eta\left(\ell-\ell^{*}\right)^{2},$ (31)
+where we have defined $\eta$ as the second derivative of the barrier with respect to $\ell$ and evaluated at $\ell^{*}$. Explicitly, this is solved for as:
+$\eta=\frac{1}{32\pi\left(\rho^{\rm in}\right)^{2}}\frac{\mathcal{V}^{4}}{\upgamma_{\rm ost}^{3}}.$ (32)
+We now consider the integral from Eq.
+(26).
+Substitution of Eq.
+(31) into Eq.
+(26), changing integration variables from $\ell$ to $\delta\ell\equiv\ell-\ell^{*}$, and performing a saddle-point approximation at $\delta\ell=0$ results in:
+$\mathcal{R}=\rho^{\rm sat}\beta_{\ell^{*}}4\pi\left(\frac{3\ell^{*}}{4\pi\rho^{\rm in}}\right)^{2/3}\sqrt{\frac{\eta}{2\pi k_{B}T^{\rm act}}}\text{exp}\left[-\frac{U(\ell^{*})}{k_{B}T^{\rm act}}\right].$ (33)
+Everything in Eq.
+(33) can now be evaluated except for the rate of single particle attachment on the critical nucleus $\beta_{\ell^{*}}$, for which we use the simple kinetic model proposed Redner [8, 2].
+$\beta_{\ell}$ is the rate of absorption of particles onto the surface of a cluster.
+This rate should be given by the density of the gas phase times the speed at which the particles swim times the probability that a given particle will absorb onto the surface:
+$\beta_{\ell}=\rho^{\rm sat}U_{o}P^{\rm abs},$ (34)
+where $P^{\rm abs}$ is the probability a particle absorbs onto the surface.
+Assuming that the surface of the nucleus has an inward pointing normal vector $\mathbf{n}$, we model $P^{\rm abs}$ as zero if a particle points away from the interface and proportional to $\mathbf{n}\cdot\mathbf{q}$ if pointing in the hemisphere towards interface.
+In other words $P^{\rm abs}$, is given by:
+$P^{\rm abs}=\begin{cases}\mathbf{n}\cdot\mathbf{q}&\mathbf{n}\cdot\mathbf{q}>0\\ 0&\mathbf{n}\cdot\mathbf{q}<0\end{cases}$ (35)
+We pick a point on the sphere with unit normal vector $-\mathbf{e}_{x}$, and then represent the orientation of a given particle as:
+$\mathbf{q}=\sin{\theta}\cos{\phi}\mathbf{e}_{x}+\sin{\theta}\sin{\phi}\mathbf{e}_{y}+\cos{\theta}\mathbf{e}_{z}.$ (36)
+Then, assuming that the orientations of the relevant particles are uniformly distributed, the average of $\beta_{\ell}$ can be expressed as:
+$\langle\beta_{\ell}\rangle=\rho^{\rm sat}U_{o}\frac{1}{4\pi}\int_{0}^{\rm 2\pi}\int_{0}^{\pi}\sin{\theta}P^{\rm abs}\text{d}\theta\text{d}\phi.$ (37)
+$P^{\rm abs}$ is only nonzero when $\phi$ is between 00 and $\pi/2$ as well as when $\phi$ is between $23\pi/$ 2 and $2\pi$, for all values of $\theta$. Then Eq. (37) reduces to:
+$\langle\beta_{\ell}\rangle=$ $\rho^{\rm sat}U_{o}\frac{1}{2\pi^{2}}\Biggl{[}\int_{0}^{\pi/2}\int_{0}^{\pi}\sin^{2}{\theta}\cos{\phi}\text{d}\theta\text{d}\phi$
+$+\int_{3\pi/2}^{2\pi}\int_{0}^{\pi}\sin^{2}{\theta}\cos{\phi}\text{d}\theta\text{d}\phi\Biggr{]}$
+$=$ $\frac{\rho^{\rm sat}U_{o}}{4}.$ (38)
+Substitution of Eq. (38) into Eq. (33) results in:
+$\mathcal{R}=\left(\rho^{\rm sat}\right)^{2}\pi U_{o}\left(\frac{3\ell^{*}}{4\pi\rho^{\rm in}}\right)^{2/3}\sqrt{\frac{\eta}{2\pi k_{B}T^{\rm act}}}\text{exp}\left[-\frac{U(\ell^{*})}{k_{B}T^{\rm act}}\right].$ (39)
+The expected time to observe a nucleation event is given by $\tau_{\rm nuc}=(\mathcal{R}V)^{-1}$, where $V$ is the total volume of the system.
+Following straightforward manipulations, it can be shown that $\eta$ is alternatively expressed as:
+$\eta=\frac{\upgamma_{\rm ost}}{2\pi(\rho^{\rm in})^{2}R_{C}^{4}}.$ (40)
+Substitution of Eq. (40) into Eq. (39) results in:
+$\mathcal{R}=\left(\rho^{\rm sat}\right)^{2}\pi U_{o}\sqrt{\frac{\upgamma_{\rm ost}}{2\pi(\rho^{\rm in})^{2}k_{B}T^{\rm act}}}\text{exp}\left[-\frac{U(\ell^{*})}{k_{B}T^{\rm act}}\right].$ (41)
+Appendix B Active Lifshitz-Slyozov-Wagner Theory
+We consider the growth of many nuclei in a supersaturated solution.
+However, we will also assume that the saturation is sufficiently small that the overall volume fraction of the final phase is effectively zero.
+In this limit the nuclei will evolve independently of one another, and we may use the single nucleus Langevin equation [Eq.
+(11a)] as the dynamics for all nuclei.
+For convenience, we will rewrite Eq.
+(11a) as:
+$\frac{\partial R}{\partial t}=\frac{1}{\zeta_{\rm eff}}F^{D}(R)+g(R)\eta(t),$ (42)
+where $\eta(t)$ is a unit Gaussian white noise with zero mean and variance $\delta(t-t^{\prime})$ and $g(R)$ is defined as:
+$g(R)=\sqrt{\frac{2k_{B}T^{\rm act}\rho^{\rm sat}}{4\pi\zeta\Delta\rho^{2}R^{3}}}.$ (43)
+and $F^{D}(R)$ remains defined as:
+$F^{D}(R)=8\pi R^{2}\upgamma_{\rm ost}\left[\frac{1}{R_{C}}-\frac{1}{R}\right].$ (44)
+Although Eq.
+(11a) contains multiplicative noise, the choice of time discretization scheme (e.g.
+Itô, Stratonovich, etc.) is inconsequential as the drift term [50] $\frac{1}{2}g(R)g^{\prime}(R)$ is $\mathcal{O}\left(R^{-4}\right)$.
+We now wish to solve for how the distribution of particle sizes evolves with time.
+This can be accomplished by first determining the Fokker-Planck equation associated with Eq.
+(42):
+$\frac{\partial P}{\partial t}=-\frac{\partial}{\partial R}\left[\frac{1}{\zeta_{\rm eff}}F^{D}(R)P(R)+\frac{1}{2}\frac{\partial}{\partial R}\left(g^{2}(R)P(R)\right)\right].$ (45)
+From Eq. (43), we can see that the diffusive term will be negligible in comparison to the drift term for large $R$, resulting in:
+$\frac{\partial P}{\partial t}=-\frac{\partial}{\partial R}\left[\frac{1}{\zeta_{\rm eff}}F^{D}(R)P(R;t)\right].$ (46)
+We can now proceed in the same fashion as Lifshitz and Slyozov [11], where it was assumed that, at long times, the asymptotic form of the distribution takes the following form:
+$P(R;t)=\frac{1}{R_{C}(t)}\mathcal{N}\left(x\right),$ (47)
+where we have defined the dimensionless radius $x\equiv R/R_{C}$ and $\mathcal{N}(x)$ is the dimensionless distribution of radii scaled by $R_{C}$.
+The time-dependence of $P(R;t)$ comes from that of the critical radius.
+Substitution of Eq.
+(47) into Eq.
+(46) results in:
+$\frac{1}{R_{C}^{2}}\frac{\partial R_{C}}{\partial t}\left[\mathcal{N}(x)+x\frac{d\mathcal{N}}{dx}\right]$
+$=\frac{2\rho^{\rm sat}\upgamma_{\rm ost}}{\zeta R_{C}^{4}\Delta\rho^{2}}\left[\frac{d\mathcal{N}}{dx}\left(x^{-1}-x^{-2}\right)+\mathcal{N}(x)\left(2x^{-3}-x^{-2}\right)\right]$ (48)
+In order for the ansatz Eq. (47) to be consistent with Eq. (48) (i.e., $\mathcal{N}(x)$ only depends on x), we require:
+$R_{C}^{2}\frac{\partial R_{C}}{\partial t}=\frac{2\rho^{\rm sat}}{\zeta\Delta\rho^{2}}\upgamma_{\rm ost}\mathcal{C},$ (49)
+where $\mathcal{C}$ is some constant. This implies that the critical radius evolves in time according to:
+$R_{C}(t)=\left[R^{3}_{C}(t_{o})+\frac{6\rho^{\rm sat}}{\zeta\Delta\rho^{2}}\upgamma_{\rm ost}\mathcal{C}t\right]^{1/3}.$ (50)
+[…]
+Appendix C Coarsening Beyond the Zero Saturation Limit
+We consider a system of phase separated active particles where the nucleation of minority phase has completed and the the final phase composition has been reached.
+In general, the system may contain many domains of minority phase.
+The overall volume fraction of this system will be finite and thus nuclei may interact with one another.
+When $\upgamma_{\rm ost}>0$, we expect the forward Ostwald ripening process to occur and large domains of minority phase to grow at the expense of small domains until the system reaches a final steady state with a single domain.
+Rather than fixing the system to be finite size and considering the influence of boundaries, we consider an infinite tiling of periodic cells.
+Due to the large system sizes and long timescales necessary to probe the dynamics of this process which evolves towards larger and larger domains, direct integration of Eq.
+(9) may be impractical.
+The problem may be greatly simplified by noting that majority phase will relax on timescales much faster than the domains of minority phase grow or shrink.
+Ignoring stochastic contributions, the density field dynamics (9) take the form:
+$\frac{\partial\rho}{\partial t}=-\frac{1}{\zeta}\bm{\nabla}\cdot\bm{\nabla}\cdot\bm{\Sigma}.$ (51)
+By restricting our perspective to only describe the majority phase, we may drop the time dependence from Eq.
+(51) by taking the quasi-static limit.
+From the perspective of the majority phase, the particles injected/adsorbed from the nuclei can be represented with finite-flux (and moving) boundary conditions at the interface between nuclei and the parent phase.
+Here, we take another approach and approximate majority phase particle addition/removal through spatially varying volumetric sources and sinks.
+This approach results in the quasi-static continuity equation now taking the following form:
+$0=-\bm{\nabla}\cdot\bm{\nabla}\cdot\bm{\Sigma}+\zeta\sum_{u,v,w}\sum_{i=1}^{N}S_{i}(\mathbf{r}),$ (52)
+where we gain the source or sink term $S_{i}$ in lieu of explicit consideration of the $i$th minority domain.
+Here the sum over $(u,v,w)$ is a summation over all periodic translations of the cell.
+Additionally, coarse-graining the effect of the minority phase on the density field into sources/sinks renders the interface between the nuclei and parent phase beyond our level of description.
+We may therefore disregard the higher order gradient terms from the dynamic stress tensor resulting in:
+$0=\bm{\nabla}^{2}\mathcal{P}(\mathbf{r})+\zeta\sum_{u,v,w}\sum_{i=1}^{N}S_{i}(\mathbf{r}).$ (53)
+Globally, the number of particles must remain conserved, imposing the following constraint on $S_{i}$:
+$\sum_{i=1}^{N}\int_{V}S_{i}(\mathbf{r})\text{d}\mathbf{r}=0.$ (54)
+From Eq.
+(53), a mathematical analogue between the pressure field and the electrostatic potential due to $N$ charge densities may be drawn.
+By approximating the minority domains as spherically symmetric we may safely collapse the “charge densities” $S_{i}$ into point charges at their centers by defining $S_{i}(\mathbf{r})\equiv 2\pi B_{i}\delta(\mathbf{r}-\mathbf{r}_{i})/\zeta$.
+Equation (53) can now be expressed as:
+$0=\bm{\nabla}^{2}\mathcal{P}(\mathbf{r})+4\pi\sum_{u,v,w}\sum_{i=1}^{N}B_{i}\delta(\mathbf{r}-\mathbf{r}_{i}-u\mathbf{a}-v\mathbf{b}-w\mathbf{c}),$ (55)
+where $\mathbf{a}$, $\mathbf{b}$, $\mathbf{c}$ are the lattice vectors and the particle conservation constraint [Eq. (54)] becomes $\sum B_{i}=0$. The solution to Eq. (55) is well-known and given by:
+$\mathcal{P}(\mathbf{r})=\mathcal{P}_{o}+\sum_{u,v,w}\sum_{i=1}^{N}\frac{B_{i}}{|\mathbf{r}-\mathbf{r}_{i}-u\mathbf{a}-v\mathbf{b}-w\mathbf{c}|},$ (56)
+where $\mathcal{P}_{o}$ is the unknown pressure of the majority phase far from the nuclei.
+We now assume that every individual nucleus is in mechanical equilibrium with its immediate surroundings.
+Eqs.
+(6) allow us to solve for the coexistence densities of a finite sized nucleus with radius $R$.
+We define $\rho^{R}(R)$ as a function mapping the size of a nucleus with size $R$ to the outside coexistence density.
+Then, if $\mathbf{r}^{s}_{j}$ is a point on the surface of a nucleus of size $R_{j}$, we have:
+$\mathcal{P}(\mathbf{r}^{s}_{j})=\mathcal{P}(\rho^{R}(R_{j}))$ (57)
+Substitution of the boundary condition [Eq. (57)] into Eq. (56) gives:
+$\mathcal{P}(\mathbf{r}^{s}_{j})$ $=\mathcal{P}(\rho^{R}(R_{j}))$
+$=\mathcal{P}_{o}+\frac{B_{j}}{R_{j}}+\sum_{u,v,w}\sum_{i=1}^{N}\frac{B_{i}}{|\mathbf{r}^{s}_{j}-\mathbf{r}_{i}-u\mathbf{a}-v\mathbf{b}-w\mathbf{c}|},$ (58)
+where the sum ignores the $i=j$ term when $u,v,w=0$.
+A key subtlety of Eq.
+(58) is that $\mathbf{r}^{s}_{j}$ is a point on the surface of nucleus $j$ while $\mathbf{r}_{i}$ is the center of nucleus $i$.
+By expanding the summation in terms of Legendre polynomials and averaging over all possible choices of $\mathbf{r}^{s}_{j}$ on the surface of the $j$th nucleus, one may show that
+$\mathcal{P}(\rho^{R}(R_{j}))=\mathcal{P}_{o}+\frac{B_{j}}{R_{j}}+\sum_{u,v,w}\sum_{i=1}^{N}\frac{B_{i}}{|\mathbf{r}_{j}-\mathbf{r}_{i}-u\mathbf{a}-v\mathbf{b}-w\mathbf{c}|},$ (59)
+where $\mathbf{r}_{j}$ is the center of nucleus $j$.
+Evaluation of the summation over an infinite periodic lattice can be accomplished via standard techniques of Ewald summation, although with the added complication that the sum ignores the $i=j$ case when $u,v,w=0$, resulting in:
+$\mathcal{P}(\rho^{R}(R_{j}))=$ $\mathcal{P}(\rho^{\rm mat})+B_{j}\left(\frac{1}{R_{j}}-\sqrt{\frac{1}{\eta\pi}}\right)+\frac{4\pi}{\mathbf{a}\cdot\left(\mathbf{b}\times\mathbf{c}\right)}\sum_{\mathbf{k}\neq\bm{0}}\sum_{i=1}^{N}\frac{B_{i}}{|\mathbf{k}|^{2}}\text{exp}\left[-\eta|\mathbf{k}|^{2}-i\mathbf{k}\cdot\left(\mathbf{r}_{j}-\mathbf{r}_{i}\right)\right]$
+$+\sum_{u,v,w}\sum_{i=1}^{N}\frac{B_{i}}{|\mathbf{r}_{j}-\mathbf{r}_{i}-u\mathbf{a}-v\mathbf{b}-w\mathbf{c}|}\text{erfc}\left(\frac{|\mathbf{r}_{j}-\mathbf{r}_{i}-u\mathbf{a}-v\mathbf{b}-w\mathbf{c}|}{2\sqrt{\eta}}\right),$ (60)
+where $\mathbf{k}$ are the reciprocal lattice vectors corresponding to our periodic cell and $\eta$ is free parameter that controls how quickly the summation converges in real space as opposed to reciprocal space.
+Equation (60) can be expressed as a matrix equation $\mathbf{b}=\mathbf{A}\cdot\mathbf{x}$, where
+$\mathbf{b}=\begin{bmatrix}\mathcal{P}(\rho^{R}(R_{1}))\\ \mathcal{P}(\rho^{R}(R_{2}))\\ \vdots\\ \mathcal{P}(\rho^{R}(R_{N}))\\ 0\end{bmatrix},$ (61a)
+$\mathbf{x}=\begin{bmatrix}B_{1}\\ B_{2}\\ \vdots\\ B_{N}\\ \mathcal{P}(\rho^{\rm mat})\end{bmatrix},$ (61b)
+and
+$\mathbf{A}=\begin{bmatrix}\frac{1}{R_{1}}+D&\mathcal{O}_{12}&\mathcal{O}_{13}&\ldots&1\\ \mathcal{O}_{21}&\frac{1}{R_{2}}+D&\mathcal{O}_{23}&\ldots&1\\ \mathcal{O}_{31}&\mathcal{O}_{32}&\frac{1}{R_{3}}+D&\ldots&1\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\ldots&0\end{bmatrix}.$ (61c)
+Here $D$ is given by:
+$D=$ $-\sqrt{\frac{1}{\eta\pi}}+\frac{4\pi}{L_{x}L_{y}L_{z}}\sum_{\mathbf{k}\neq 0}\frac{1}{|\mathbf{k}|^{2}}\text{exp}\left(-\eta|\mathbf{k}|^{2}\right)$
+$+\sum_{u,v,w\neq 0}\frac{1}{|\mathbf{r}_{uvw}|}\text{erfc}\left(\frac{|\mathbf{r}_{uvw}|}{2\sqrt{\eta}}\right),$ (62)
+where we have defined ${|\mathbf{r}_{uvw}|\equiv\left(|u\mathbf{a}|^{2}+|v\mathbf{b}|^{2}+|w\mathbf{c}|^{2}\right)^{1/2}}$ and $\mathcal{O}_{ij}$ is given by:
+$\mathcal{O}_{ij}=$ $\frac{4\pi}{\mathbf{a}\cdot\left(\mathbf{b}\times\mathbf{c}\right)}\sum_{\mathbf{k}\neq\bm{0}}\frac{1}{|\mathbf{k}|^{2}}\text{exp}\left(-\eta|\mathbf{k}|^{2}-i\mathbf{k}\cdot\left(\mathbf{r}_{j}-\mathbf{r}_{i}\right)\right)$
+$+\sum_{u,v,w}\frac{1}{|\mathbf{r}_{uvw}^{ij}|}\text{erfc}\left(\frac{|\mathbf{r}_{uvw}^{ij}|}{2\sqrt{\eta}}\right),$ (63)
+where we have defined ${|\mathbf{r}^{ij}_{uvw}|\equiv|\mathbf{r}_{j}-\mathbf{r}_{i}-u\mathbf{a}-v\mathbf{b}-w\mathbf{c}|}$.
+Given the size and locations of each nucleus in the cell, one may then solve Eqs.
+(61) in order to find all $B_{i}$ and the background pressure $\mathcal{P}_{o}$.
+Let’s now consider the rate of change of the radius of a particular nucleus and how it translates to the nucleus source/sink strength $B_{i}$.
+If nucleus $i$ changes its radius $R_{i}$ by an amount $dR_{i}$ over a time period $dt$, then the rate of mass injected into the majority phase by particle $i$ is given by:
+$4\pi R_{i}^{2}\frac{dR_{i}}{dt}\left(\rho^{\rm out}-\rho^{\rm in}\right)$ (64)
+By the definition of $S_{i}$, the rate of mass injected into the majority phase by nucleus $i$ is also equal to the integral of $S_{i}$ across the entire system of volume. Then we have:
+[…]
+From the definition of $B_{i}$ we then have:
+$4\pi R_{i}^{2}\frac{dR_{i}}{dt}\left(\rho^{\rm out}-\rho^{\rm in}\right)=\frac{4\pi}{\zeta}B_{i}$ (66)
+Then the rate of change of the radius of nucleus $i$ is given by:
+$\frac{dR_{i}}{dt}=\frac{B_{i}}{\zeta R_{i}^{2}(\rho^{\rm out}-\rho^{\rm in})}$ (67)
+Therefore solution of the matrix equation defined by Eqs. (61) allows for the extraction of the coarsening rates.